### v1

# Schur’s Positive-Definite Network: Deep Learning in the SPD cone with structure

Abstract
Estimating matrices in the symmetric positive-definite (SPD) cone is of interest for many applications ranging from computer vision to graph learning. While there exist various convex optimization-based estimators, they remain limited in expressivity due to their model-based approach. The success of deep learning has thus led many to use neural networks to learn to estimate SPD matrices in a data-driven fashion. For learning structured outputs, one promising strategy involves architectures designed by unrolling iterative algorithms, which potentially benefit from inductive bias properties. However, designing correct unrolled architectures for SPD learning is difficult: they either do not guarantee that their output has all the desired properties, rely on heavy computations, or are overly restrained to specific matrices which hinders their expressivity. In this paper, we propose a novel and generic learning module with guaranteed SPD outputs called SpodNet, that also enables learning a larger class of functions than existing approaches. Notably, it solves the challenging task of learning jointly SPD and sparse matrices. Our experiments demonstrate the versatility of SpodNet layers.
1 Introduction
The estimation of symmetric positive-definite (SPD) matrices is a major area of research, due to their crucial role in various fields ranging such as optimal transport (Bonet et al., 2023; Ju and Guan, 2022; Yair et al., 2019), graph theory (Lauritzen, 1996), computer vision (Nguyen et al., 2019) or finance (Ledoit and Wolf, 2003). While various statistical estimators, i.e. model-based, have been developed (Ledoit and Wolf, 2004; Banerjee et al., 2008; Cai et al., 2011), a more recent trend is to use learning-based approaches to estimate SPD matrices with neural networks (Gao et al., 2020).
Training neural networks while enforcing non-trivial structural constraints such as positive-definiteness is a difficult task. There have been many efforts in this direction in recent years, often in an ad-hoc manner and each with their own shortcomings (see Section 2 for more details). Building on the seminal work of Gregor and LeCun (2010) in sparse coding, a promising research direction involves designing neural networks architectures from the unrolling of an optimization algorithm (Chen and Pock, 2016; Monga et al., 2021; Chen et al., 2022). In the case of SPD matrices, algorithm unrolling presents several challenges. First, algorithms operating in the SPD cone usually rely on heavy operations such as retractions (Boumal, 2023), SVD or line search (Rolfs et al., 2012). These operations do not integrate well into a neural network architecture, making these algorithms difficult to unroll.
Second, beyond ensuring positive definiteness, it is often desirable to impose additional structure on the matrix to estimate, such as elementwise sparsity. Simultaneously enforcing SPD + sparsity constraints on the output of a neural architecture is extremely difficult (Guillot and Rajaratnam, 2015; Sivalingam, 2015). So far, it is only a hypothetical property that is not satisfied in practice (see Figure 2 and Section 5 for more details).
Contributions: In this paper, we combine the strengths of both model-based and data-driven approaches and we make the following contributions:
  • We introduce a novel framework for building learning modules with a guaranteed SPD-to-SPD mapping for which it is also possible to enforce additional constraints on the output. We refer to it as SpodNet for Schur’s Positive-Definite Network. Our approach is originally inspired by the block-matrix updates perspective of the Graphical Lasso solver (Mazumder and Hastie, 2012) and extended to generic SPD learning. As a particular case, SpodNet layers are able to learn jointly SPD and sparse matrices in a data-driven fashion. To the best of our knowledge, SpodNet-based architectures are the first to offer strict guarantees combining both properties.
  • We illustrate the framework’s versatility through applications in sparse precision matrix estimation. In this scenario, we present limitations of other learning-based approaches and demonstrate how to overcome them with SpodNet layers. Our experiments validate the use of SpodNet for learning jointly SPD-to-SPD and sparsity inducing functions, and show promising results across various performance metrics.
Notation
We reserve the bold uppercase for matrices $\bm{\Theta}$, bold lowercase $\bm{\theta}$ for vectors and standard lowercase $\theta$ for scalars. The soft-thresholding function is $\operatorname{ST}_{\gamma}(\cdot)=\operatorname{sign}(\cdot)\max(\lvert\cdot\rvert-\gamma,0)$ for $\gamma\geq 0$; it acts elementwise on vectors or matrices. On matrices, $\lVert\cdot\rVert_{1}$ is the sum of absolute values of the matrix coefficients. The cone of SPD matrices in dimension $p$ is denoted $\mathbb{S}^{p}_{++}$. We use $\bm{\Sigma}$ to designate covariance matrices, $\bm{\Theta}$ for precision matrices, and $\bm{S}$ for empirical covariance matrices.
2 Related works
Before delving into the details of SpodNet, we briefly review existing approaches for constructing SPD matrices, which can be broadly divided into three categories.
Riemannian learning
Riemannian optimization provides tools to build algorithms whose iterates lie on Riemannian manifolds, such as the SPD manifold (Absil et al., 2008). Many have adopted these tools to design neural architectures that operate on the manifold of SPD matrices (Huang and Van Gool, 2017; Gao et al., 2020, 2022). However, a common and major bottleneck of these models is their dependence on Riemannian operations which are notoriously costly to compute. Furthermore, Riemannian optimization beyond smooth objective functions is still an active area of research (Absil and Hosseini, 2019; Huang and Wei, 2022; Zhou et al., 2023) and it is thus highly non-trivial to design neural architectures to handle additional constraints such as elementwise sparsity.
Projection-based methods
For some applications, it is of interest to project the matrices from a high-dimensional SPD manifold to a lower-dimensional one (Dong et al., 2017; Nguyen et al., 2019). Using linear algebra principles, these approaches design specific layers to guarantee that the final matrix is SPD. However, these solutions are ad-hoc and lack straightforward extensions for general SPD-to-SPD learning with additional structural constraints. Specifically, there is no obvious way to incorporate elementwise sparsity on the matrices without risking breaking their SPD structure. For instance, hard-thresholding the off-diagonal entries of a SPD matrix does not always preserve the SPD property (Guillot and Rajaratnam, 2012) and, more generally, elementwise functions preserving this property are very limited (Guillot and Rajaratnam, 2015).
Unrolled learning
In order to ensure that a network’s output strictly respects some desired properties, another recent and successful direction is to unroll convex optimization algorithms (Chen et al., 2022). This unrolling procedure acts as an “inductive bias” on the architecture (Shrivastava, 2020), naturally guiding the model to explore suitable solutions within the space of imposed properties. In SPD learning, this approach has been exploited by Shrivastava et al. (2019) who unrolled an optimization algorithm to train neural networks to estimate sparse inverses of covariance matrices. Whilst algorithm unrolling is a very powerful approach to learn specific matrices, it is very difficult to extend it to generic SPD learning. We provide more details about the limits of related work in this setting in Section 4.2.
3 The SpodNet framework
We now introduce the SpodNet layer which is at the core of our contribution. Essentially, it is an SPD-to-SPD mapping parameterized by neural networks. As described in Section 4, SpodNet is inspired by an algorithm used to solve the Graphical Lasso in Mazumder and Hastie (2012).
A SpodNet layer operates by cycling through the $p$ column-row pairs individually and i) updating the corresponding column-row with a neural network, and ii) updating the diagonal element to maintain the SPD constraint (see Figure 1). Interestingly, the neural network at step i) can update the column-row to any value without compromising the SPD guarantee. Consequently, one is free to exploit a spectrum of approaches for those updates depending on other desired structural properties of the output matrix. In particular, we describe in Section 4 a specific implementation of SpodNet that enforces elementwise sparsity for learning sparse precision matrices.
1:  Input: $\bm{\Theta}_{\mathrm{in}}\in\mathbb{S}_{++}^{p}$ and $\bm{W}_{\mathrm{in}}=\bm{\Theta}_{\mathrm{in}}^{-1}$
2:  for column $i\in\{1,\cdots,p\}$ do
3:     Extract blocks: $\bm{W}_{11}$, $\bm{w}_{12},w_{22}$
4:     Compute $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$
5:     New column $\bm{\theta}_{12}^{+}=f(\bm{\Theta})$
6:     New diagonal value $\theta_{22}^{+}=g(\bm{\Theta})+{\bm{\theta}_{12}^{+}}^{\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}$
7:     Update $\bm{\Theta}=\bm{\Theta}^{+},\bm{W}=\bm{W}^{+}$ as in Equations 2 and 3
8:  end for
9:  Output: $\bm{\Theta}_{\mathrm{out}}\in\mathbb{S}_{++}^{p}$ and $\bm{W}_{\mathrm{out}}=\bm{\Theta}_{\mathrm{out}}^{-1}$
Algorithm 1 One SpodNet layer
Figure 1: Schematic overview of SpodNet layers. Our framework chains updates of column-row pairs and diagonals using neural networks (note: the updated columns and rows are not constant, despite being represented by constant colors). At all times the matrices remain SPD via Schur’s condition.
3.1 Algorithmic foundations
More formally, when the SpodNet layer operates on the $i$-th column and row of $\bm{\Theta}\in\mathbb{S}_{++}^{p}$ it also updates the $i$-th diagonal coefficient accordingly. The key to preserving the positive-definiteness of the matrix upon changing its $i$-th column and row is an appropriate update of the diagonal entry $\Theta_{ii}$ based on Schur’s condition for positive-definiteness. In the following, a $+$ superscript on a variable (e.g. $\bm{\theta}^{+}$) indicates an update value of this variable (e.g. outputted by a neural network). The following proposition formalizes the SpodNet operations.
Proposition 3.1.
Suppose the updated column-row pair is the last one ($i=p)$. We partition $\bm{\Theta}$ as
$\bm{\Theta}=\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}_{12}\\ \bm{\theta}_{21}&\theta_{22}\end{bmatrix},\quad\text{with}\quad\bm{\Theta}_{11}\in\mathbb{R}^{(p-1)\times(p-1)},\,\bm{\theta}_{12}\in\mathbb{R}^{p-1},\ \bm{\theta}_{21}=\bm{\theta}_{12}^{\top},\,\theta_{22}\in\mathbb{R}\,.$ (1)
(for a generic column $i$, $\bm{\Theta}_{11}$ refers to $\bm{\Theta}$ without its $i$-th row and $i$-th column, $\bm{\theta}_{12}$ is the $i$-th row of $\bm{\Theta}$ without its $i$-th value, and $\theta_{22}$ is $\Theta_{ii}$ as illustrated in Algorithm 1).
Let $f:\mathbb{R}^{p\times p}\rightarrow\mathbb{R}^{p-1}$, $g:\mathbb{R}^{p\times p}\rightarrow\mathbb{R}_{++}$, where $f$ and $g$ are arbitrary functions. Then, updating the $i$-th row and column of $\bm{\Theta}$ as
$\bm{\Theta}^{+}\triangleq\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}^{+}_{12}\triangleq f(\bm{\Theta})\\ \bm{\theta}_{21}^{+}\triangleq f(\bm{\Theta})^{\top}&\theta_{22}^{+}\triangleq g(\bm{\Theta})+\bm{\theta}^{+}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}^{+}_{12}\end{bmatrix}\,,$ (2)
preserves the SPD property, i.e. $\bm{\Theta}^{+}\in\mathbb{S}_{++}^{p}$.
Proof.
Since $\bm{\Theta}$ is positive-definite, so is its leading principal submatrix $\bm{\Theta}_{11}$. It follows that $\bm{\Theta}^{+}$ is well-defined, and obviously symmetric. Its positive-definiteness then ensues from Schur’s condition for positive-definiteness. Indeed $\bm{\Theta^{+}}$ is SPD when $\bm{\Theta}_{11}\succ 0$ and $\theta_{22}^{+}-\bm{\theta}_{12}^{+\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}=g(\bm{\Theta})>0$ (Zhang, 2006, Theorem 1.12). The latter condition is ensured as the output of $g$ is a positive scalar. ∎
Finally, one SpodNet layer chains $p$ updates of the form (2), sequentially updating all column-row pairs one after the other as summarized in Algorithm 1. A full SpodNet architecture then stacks $K$ SpodNet layers. The overall procedure is schematized in Figure 1 for $K=2$.
The expressivity of SpodNet comes from the simple realization that one is free to plug any arbitrary functions $f$ and $g$ in the updates whilst guaranteeing that $\bm{\Theta}$ remains in the SPD cone. Namely, additional structure such as sparsity can trivially be imposed on $\bm{\Theta}$ through $f$. We next detail the computational cost of fully updating $\bm{\Theta}$ through one SpodNet layer.
3.2 Improving update complexity
The update of Equation 2 a priori requires inverting the matrix $\bm{\Theta}_{11}$ which comes with a prohibitive cost of $\mathcal{O}(p^{3})$ for each column-row update. We can fortunately leverage the column-row structure of the updates to decrease the cost to a mere $\mathcal{O}(p^{2})$, by maintaining the matrix $\bm{W}=\bm{\Theta}^{-1}$ up-to-date along the iterations.
Proposition 3.2.
Let $\bm{W}$ be the inverse of $\bm{\Theta}$, adopting the same block structure $\bm{W}=\begin{bmatrix}\bm{W}_{11}&\bm{w}_{12}\\ \bm{w}_{12}^{\top}&w_{22}\end{bmatrix}$. Then $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$. In addition, if $\bm{\Theta}$ is updated as $\bm{\Theta}^{+}$ following Equation 2, then the update of $\bm{W}$ defined by
$\bm{W}^{+}\triangleq\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}\bm{\theta}_{21}^{+}[{\bm{\Theta}}_{11}]^{-1}}{g(\bm{\Theta})}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}}{g(\bm{\Theta})}\\ \cdot&1/g(\bm{\Theta})\end{bmatrix}\,.$ (3)
can be computed in $\mathcal{O}(p^{2})$ and satisfies $[\bm{W}^{+}]^{-1}=\bm{\Theta}^{+}$.
Proof.
By using the Banachiewicz inversion formula on Schur’s complement (Zhang, 2006, Theorem 1.2)
$\begin{bmatrix}\bm{W}_{11}&\bm{w}_{12}\\ \bm{w}_{21}&w_{22}\end{bmatrix}$ $=\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}_{12}\\ \bm{\theta}_{21}&\theta_{22}\end{bmatrix}^{-1}=\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\\ \\ \cdot&\frac{1}{\theta_{22}-\bm{{\theta}}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\end{bmatrix}\,.$ (4)
Identifying all blocks yields $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$ which can be computed in $\mathcal{O}(p^{2})$ if one has access to $\bm{W}$. This can be achieved using Equation 3, and involves only operations in $\mathcal{O}(p^{2})$. The property $\bm{\Theta}^{+}=[\bm{W}^{+}]^{-1}$ is statisfied by using the same inversion formula. ∎
Based on this result a full update of $\bm{\Theta}$ can be achieved with cost $\mathcal{O}(p^{3})$ ($p$ column-row updates of cost $\mathcal{O}(p^{2})$).
4 Learning sparse precision matrices with SpodNet
In this section, we propose three specific SpodNet-based architectures for learning sparse precision matrices, which naturally arise in the context of Gaussian graphical models (Lauritzen, 1996) that we now briefly recall.
4.1 A model-based formulation: the Graphical Lasso
Consider a dataset of $n$ observed signals $\bm{x}_{1},\ldots,\bm{x}_{n}$ where each $\bm{x}_{i}\in\mathbb{R}^{p}$ follows a certain (centered) distribution with covariance matrix $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$. The problem of precision matrix estimation arises when attempting to identify the conditional dependency graph of the $p$ variables of this dataset. Estimating these dependencies is of interest in many fields such as neuroscience (Rubinov and Sporns, 2010), finance (Ledoit and Wolf, 2003) or genomics (Cai et al., 2013). In Gaussian graphical models, this graph is associated with the so-called precision matrix $\bm{\Theta}=\bm{\Sigma}^{-1}$ which directly encodes the dependencies [^1] In this case $\Theta_{ij}=0$ iff variable $i$ and $j$ are conditionally independent, given the other variables.. The problem of precision matrix estimation involves determining $\bm{\Theta}\in\mathbb{S}^{p}_{++}$ given the empirical covariance matrix $\bm{S}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}$. Moreover, in many practical situations the graph (thus the matrix $\bm{\Theta}$) is also sparse due to limited conditional dependencies, which leads to a sparse precision matrix estimation problem.
In this case, a very popular estimator is the Graphical Lasso (GLasso) (Banerjee et al., 2008; Friedman et al., 2008) that solves the convex optimization problem
$\min_{\bm{\Theta}\succ 0}\ -\log\det(\bm{\Theta})+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{\Theta}\|_{1,\mathrm{off}}\,,$ (5)
where $\|\bm{\Theta}\|_{1,\mathrm{off}}$ denote the off-diagonal $\ell_{1}$ norm of $\bm{\Theta}$, equal to $\sum_{i\neq j}|\Theta_{ij}|$. Equation 5 corresponds to an $\ell_{1}$-penalized maximum likelihood estimator under the assumption that the signals are Gaussian with covariance $\bm{\Sigma}$. It can be computed with various convex optimization solvers, e.g. primal formulations (Mazumder and Hastie, 2012), proximal gradient descent (Rolfs et al., 2012), or second order algorithms (Hsieh et al., 2014), all with strict guarantees of recovering both sparse and SPD precision matrices. In particular, the algorithm of Mazumder and Hastie (2012) precisely fits the form of the updates in (2), where the functions $f$ and $g$ are not learnt, but correspond to solutions of changing Lasso regression problems.
Learning-based approaches, however, are scarcer. To the best of our knowledge, DeepGraph (Belilovsky et al., 2017) and GLAD (Shrivastava et al., 2019) are the only two existing works to tackle sparse precision matrix estimation under a data-driven perspective, both with restrictive limitations. This somewhat surprising fact is justified by the challenges posed by the task, which we highlight in the sequel.
4.2 Limitations of learning-based approaches for sparse precision matrix estimation
DeepGraph.Belilovsky et al. (2017) propose to directly map empirical covariance matrices to the graph structures using a CNN-based architecture trained with supervision. DeepGraph is trained for structure recovery only and not for the estimation of the actual precision matrix. It disregards the magnitude of the elements of the output matrix and focuses solely on its support, thus circumventing the problem of SPD learning, which is a limitation for some applications.
Figure 2: GLAD limitations: Smallest eigenvalue (orange), density degree (green) and relative discrepancy of the two matrices (blue) for one of the output $(\bm{Z}_{\mathrm{out}},\bm{\Theta}_{\mathrm{out}})$ of GLAD. $\bm{\Theta}_{\mathrm{out}}$ and $\bm{Z}_{\mathrm{out}}$ are significantly different, $\bm{\Theta}_{\mathrm{out}}$ is not sparse, and $\bm{Z}_{\mathrm{out}}$ is not positive-definite.
GLAD. Shrivastava et al. (2019) unroll an alternating minimization algorithm for the Graphical Lasso (Dalal and Rajaratnam, 2017) to serve as an inductive bias for their model, which learns to predict elementwise regularization parameters of the updates. GLAD is trained with supervision, by minimizing a reconstruction error between the outputs and the ground truth matrix. The unrolled algorithm is designed to solve a Lagrangian relaxation of (5)
$\min_{\bm{\Theta},\bm{Z}\in\mathbb{S}^{p}_{++}}-\log\det\left(\bm{\Theta}\right)+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{Z}\|_{1}+\frac{\alpha}{2}\|\bm{Z}-\bm{\Theta}\|_{F}^{2}\,.$ (6)
As detailed in Section A.1, GLAD jointly operates on two matrices $\bm{Z}$ and $\bm{\Theta}$. GLAD is a powerful model, and has a minimalist yet effective parametrization. Unfortunately, by treating $\bm{Z}$ and $\bm{\Theta}$ separately, the architecture cannot learn matrices that are both sparse and SPD. This is illustrated in Figure 2 where $\bm{\Theta}$ and $\bm{Z}$ differ, $\bm{\Theta}$ is not sparse and $\bm{Z}$ is not SPD.
4.3 New models: SpodNet for learning sparse precision matrices
In this section, we present how to leverage SpodNet layers to overcome the limitations of existing solutions for learning both sparse and SPD precision matrices. By order of complexity, we present three neural architectures relying on SpodNet layers, each corresponding to a specific choice of the functions $f$ and $g$ in Equation 2.
All three architectures are inspired by an unrolling of a proximal block coordinate descent applied to the Graphical Lasso problem (5), where the blocks are column-row pairs described in Section 3. Because (full) proximal gradient descent on (5) is called Graphical ISTA (Rolfs et al., 2012), we coin this Block-Graphical ISTA.
To derive its iterations, we note that the gradient of $\bm{\Theta}\mapsto-\log\det(\bm{\Theta})+\langle\bm{S},\bm{\Theta}\rangle$ (the smooth term in the GLasso loss) is $-\bm{\Theta}^{-1}+\bm{S}$ (Boyd and Vandenberghe, 2004, Section A.4.1) and thus the gradient restricted to the $\bm{\theta}_{12}$-block is $-[\bm{\Theta}^{-1}]_{12}+\bm{s}_{12}=-\bm{w}_{12}+\bm{s}_{12}$. Consequently, the proximal gradient descent algorithm on these coordinates writes $\bm{\theta}_{12}^{+}=\mathrm{prox}_{\lambda\gamma\|\cdot\|_{1}}(\bm{\theta}_{12}-\gamma(\bm{s}_{12}-\bm{w}_{12}))$ where $\gamma>0$ is some stepsize and $\mathrm{prox}$ the proximal operator.
Using that the soft-thresholding is the proximal operator of the $\ell_{1}$ norm, Block-Graphical ISTA hence reads
$\bm{\theta}_{12}^{+}=\operatorname{ST}_{\gamma\lambda}(\bm{\theta}_{12}-\gamma(\bm{s}_{12}-\bm{w}_{12}))\,.$ (7)
Unrolled Block Graphical-ISTA (UBG)
First, we propose to learn the stepsizes and soft-thresholding levels when unrolling the Block Graphical ISTA iterations (7). In a learning-based approach, we use as updating function $f$,
$f_{\mathrm{UBG}}:\bm{\theta}_{12}\mapsto\operatorname{ST}_{\Lambda^{+}}(\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12}))\,,$ (8)
where the elementwise soft-thresholding parameters $\Lambda^{+}$ along with the step-size $\gamma^{+}$ are predicted by perceptrons at each update. We emphasize that this is different from hyperparameter tuning, since $\gamma^{+}$ and $\Lambda^{+}$ are predicted at each update instead of being global parameters. UBG exhibits the highest inductive bias among the models, as it’s architecture is directly derived from unrolling an iterative optimization algorithm designed to minimize a model-based loss. Consequently, this architecture is particularly suited for learning sparse precision matrices.
However, as highlighted in Section 3, the true power of SpodNet lies in the fact that one can plug any update functions and still get SPD outputs. This first motivates extending UBG to a more expressive setting with less inductive bias.
Plug-and-Play Block Graphical-ISTA (PNP)
Next, we extend UBG to the Plug-and-Play setting (Venkatakrishnan et al., 2013; Romano et al., 2017; Kamilov et al., 2023). In a nutshell, these type of methods replace the proximal operator in first-order algorithms by a denoiser, usually implemented by a neural network. In our context, this corresponds to replacing the soft-thresholding of UBG by an arbitrary network $\Psi:\mathbb{R}^{p-1}\to\mathbb{R}^{p-1}$, that is,
$f_{\mathrm{PnP}}:\bm{\theta}_{12}\mapsto\Psi(\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12}))\,.$ (9)
To promote sparsity of the output of $f_{\mathrm{PnP}}$, the last layer of $\Psi$ performs an elementwise soft-thresholding. The parameter for this soft-thresholding is also learned from the data and given by the same perceptron that predicts $\Lambda^{+}$ in UBG.
End-to-end Learnable Block Graphical-ISTA (E2E)
Finally, we propose a fully-flexible architecture without any algorithm-inspired assumptions. Precisely, we consider
$f_{\mathrm{E2E}}:\bm{\theta}_{12}^{+}\mapsto\Phi(\bm{\theta}_{12})\,,$ (10)
where $\Phi$ takes in the current state of the column $\bm{\theta}_{12}$ and learns to predict an adequate column update $\bm{\theta}_{12}^{+}$. Intuitively, the neural network $\Phi$ acts as learning both the forward and the backward steps of a forward-backward iteration (Combettes and Pesquet, 2011). E2E is the most expressive of our three models because it is free from any inductive bias. As for PNP, sparsity of the predictions is enforced by a soft-thresholding non-linearity in the last layer of $\Phi$.
Figure 3: Potential instabilities with no normalization: Along column update indices, we plot the smallest eigenvalue (blue), largest diagonal value (orange) and conditioning (green) of $\bm{\Theta}$. We observe that $\bm{\Theta}$ remains SPD but gets increasingly closer to being singular during the updates.
Stability and update of the diagonal coefficient $\theta_{22}$
Although the positive definiteness of $\bm{\Theta}$ is guaranteed to be preserved at each update for any positive-valued function $g$, we have empirically observed that its smallest eigenvalue could approach 00 as visible in Figure 3. In practice this leads to instability in training. Below, we offer an interpretation of this phenomenon and suggest a solution to address it. Each column-row update inside a SpodNet layer can be written as the rank-$2$ update
$\bm{\Theta}^{+}=\underbrace{\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}_{12}\\ \bm{\theta}_{21}&\theta_{22}\end{bmatrix}}_{=\bm{\Theta}}+\underbrace{\begin{bmatrix}0&\bm{\theta}_{12}^{+}-\bm{\theta}_{12}\\ (\bm{\theta}_{12}^{+}-\bm{\theta}_{12})^{\top}&\theta_{22}^{+}-\theta_{22}\end{bmatrix}}_{\triangleq\Delta_{\bm{\Theta}}}\,.$ (11)
Moreover, the nonzero eigenvalues of the perturbation $\Delta_{\bm{\Theta}}$ are given by $\lambda_{\pm}=\frac{(\theta_{22}^{+}-\theta_{22})\pm\sqrt{(\theta_{22}^{+}-\theta_{22})^{2}+4\|\bm{\theta}_{12}^{+}-\bm{\theta}_{12}\|^{2}}}{2}$, and by the Bauer-Fike theorem, we can quantify the evolution of $\bm{\Theta}$’s spectrum by
$|\lambda_{k}(\bm{\Theta})-\lambda_{k}(\bm{\Theta}^{+})|\leq\|\Delta_{\bm{\Theta}}\|_{\mathrm{op}}\,,$ (12)
where $\lambda_{k}$ denotes the $k$-th largest eigenvalue of a SPD matrix. Hence, one way to ensure that the perturbation of the spectrum is small is to control $\|\Delta_{\bm{\Theta}}\|_{\mathrm{op}}$, for instance, by limiting the magnitude of the updates. Experimentally, the most successful approach to control $\|\Delta_{\bm{\Theta}}\|_{\mathrm{op}}$ is to limit the magnitude of $\theta^{+}_{22}$, which is not a trivial task as the updated value $\theta^{+}_{22}=g(\bm{\Theta})+\bm{\theta}^{+}_{21}[\bm{\Theta}_{11}]^{-1}\bm{\theta}^{+}_{12}$ must satisfy $\theta^{+}_{22}-\bm{\theta}^{+}_{21}[\bm{\Theta}_{11}]^{-1}\bm{\theta}^{+}_{12}>0$. Thus, in practice we propose to scale $\bm{\theta}_{12}^{+}$, ultimately choosing for all 3 models $f\in\{f_{\mathrm{UBG}},f_{\mathrm{PnP}},f_{\mathrm{E2E}}\}$ the column update
$\bm{\theta}_{12}^{+}=\sqrt{\zeta}\frac{f(\bm{\Theta})}{\sqrt{f(\bm{\Theta})^{\top}[\bm{\Theta}_{11}]^{-1}f(\bm{\Theta})}}\,,$ (13)
which ensures that the diagonal update is given by $\theta_{22}^{+}=g(\bm{\Theta})+\zeta$. The hyperparameter $\zeta>0$ acts as a form of regularization that handles a compromise between stability and expressivity of the model. In all of our experiments we set $\zeta=1$.
Finally, for all models, we use for $g$ a small neural network that takes as input the previous value of $\theta_{22}$ and the corresponding diagonal entry of $\bm{S}$, $s_{22}$, and whose positivity is ensured by using an exponential function as final nonlinearity.
5 Experiments
We now illustrate our three graph learning models built on SpodNet layers’s ability to learn relevant sparse precision matrices. This section aims to achieve two objectives. First, we highlight that our three models learn matrices that are indeed sparse and SPD in contrast to other learning-based approaches such as GLAD. Second, we show that our models identify accurate sparse precision matrices and even outperform an oracle Graphical Lasso estimator.
Figure 4: Training dynamics of our UBG, PNP and E2E models, compared to the baseline Oracle GLasso (in red) which provides both sparse and SPD matrices.
Data
We use synthetically generated data to fit our models. We generate $N$ sparse SPD matrices of dimension $p=50$ using sklearn’s make_sparse_spd_matrix function (Pedregosa et al., 2011), of which we ensure proper conditioning by adding $0.1$ to the diagonal entries. These matrices are treated as ground truth precision matrices $\bm{\Theta}_{\mathrm{true}}^{(i)}$ for $i\in[N]$. The sparsity degree of each matrix is controlled through the one imposed on their Cholesky factors during their generation. We consider two setups for these sparse SPD matrices: weakly sparse ones ($\alpha=0.7$), and strongly sparse ones ($\alpha=0.95$). From each of these $N$ ground truth precision matrices $\bm{\Theta}_{\mathrm{true}}=\bm{\Sigma}_{\mathrm{true}}^{-1}$, we sample $n=50$ i.i.d. centered Gaussian random vectors $\bm{x}_{j}\sim\mathcal{N}(0,\bm{\Sigma}^{(i)}_{\mathrm{true}}),j\in[n]$, which are used to compute an empirical covariance matrix $\bm{S}^{(i)}$. Both the training and testing datasets thus comprise distinct couples $(\bm{S}^{(i)},\bm{\Theta}^{(i)}_{\mathrm{true}})$, meaning that each $\bm{S}^{(i)}$ stems from a different distribution. Since $n=p$, the $\bm{S}^{(i)}$’s are poorly conditioned and the estimation of $\bm{\Theta}_{\mathrm{true}}^{(i)}$ is typically considered to be a very difficult problem to solve. We generate $N_{\mathrm{test}}=100$ $(\bm{\Theta},\bm{S})$ couples for the testing set on which we evaluate all approaches. The average sparsity degrees of the ground truth in the weakly sparse and strongly sparse testing sets are respectively about 24% and 90%. The training set for our UBG model comprises 100 couples, whereas PNP and E2E are trained on 1000 couples.
Minimization objective & performance metric
All three models are trained using ADAM with default parameters (Kingma and Ba, 2014) to minimize a reconstruction error loss $\mathcal{L}_{\mathrm{MSE}}=\frac{1}{N_{\mathrm{train}}}\sum^{N_{\mathrm{train}}}_{i=1}\|\hat{\bm{\Theta}}^{(i)}-\bm{\Theta}^{(i)}_{\mathrm{true}}\|^{2}_{F}$ in the vein of GLAD, where $N_{\mathrm{train}}$ is the size of the training set and $\hat{\bm{\Theta}}$ is the model’s output.
Architectures & training
We consider in our experiments models with $K=1$ SpodNet layer. The input to the models are the empirical covariance matrices $\bm{S}$, from which we compute $\bm{\Theta}_{\mathrm{init}}=(\bm{S}+0.1*\bm{I})^{-1}$ and $\bm{W}_{\mathrm{init}}=\bm{S}+0.1*\bm{I}$, which serve as the inputs to the SpodNet layer. We use learning rates of $0.01$, $0.02$ and $0.03$ respectively, to train UBG, PNP and E2E. The principal difficulty in training UBG and PNP stems from learning to predict $\gamma^{+}$. We find that using a perceptron with a single hidden layer of 10 neurons and ReLU activations that takes in the current state of the whole matrix to predict a $\gamma^{+}$ is expressive enough to learn adequate matrices (Figures 4 and 6). Its positivity is ensured by taking the absolute value of the prediction. We believe that a more extensive architectural search for this perceptron could help achieve even better results. The prediction of the $\Lambda^{+}$ parameters at each step is performed by a small single layer perceptron of 5 neurons with ReLU activation functions and an absolute value function following the last layer to ensure positivity. This network acts elementwise on the entries of the input vector. The sufficiency of a lower amount of training data for UBG could be imputed to its stronger inductive bias. PNP’s $\Psi$ function and E2E’s $\Phi$ function are learned by a single layer ReLU perceptron of 500 neurons. All experiments were ran on an RTX 2080 SUPER GPU with an AMD EPYC 7302 CPU and resources of similar computing power.
Sparsity & positive-definiteness
Figure 5 highlights that all three of our models are indeed able to learn jointly sparse and SPD matrices. Contrary to GLAD, the raw predictions respect these structural constraints by construction. For the sake of completeness, we threshold the entries of GLAD’s raw predictions at increasing values to artificially introduce sparsity, and observe that this can indeed break their positive-definiteness. As visual confirmation, Figure 6 shows one of UBG and PNP’s predictions in the weakly sparse and strongly sparse settings.
Performance & benchmark
Since GLAD’s predictions are not naturally sparse and SPD, we benchmark against sklearn’s implementation of the GLasso solver (Friedman et al., 2008), which is the standard algorithm for learning sparse precision matrices. We compare ourselves to an ideal Oracle GLasso where the $\ell_{1}$ hyperparameter $\lambda^{(i)}$ is individually assigned to the solver for each target matrix $\bm{\Theta}_{\mathrm{true}}^{(i)}\forall i\in[N_{\mathrm{test}}]$ and where $\lambda^{(i)}$ is optimally tuned so as to achieve best possible reconstruction error. In practice, we validate this $\lambda^{(i)}$ in $[10^{-3}\lambda_{\max},\lambda_{\max}]$ where $\lambda_{\max}$ is the parameter such that every off-diagonal entry of the prediction is zero (see Section A.2). We emphasize that this is an idealistic scenario where the GLasso is at its most favorable. The results are depicted in Figure 4 for the weakly sparse setting and illustrates that our three models are indeed able to learn jointly sparse and SPD matrices while outperforming the Oracle GLasso solver. PNP yields the best performance in terms of NMSE whereas the E2E model is best in terms of support recovery in our setting. Figure 6 offers visual confirmation of the relevancy of the predictions. Interestingly, merely only one SpodNet layer is enough to achieve good performances for all of the three models.
Figure 5: Sparsity (left) and positive-definiteness (right) of the raw predictions of all three of our models and GLAD’s, along with the latter’s thresholded predictions, on the testing set of the weakly sparse setting.
Figure 6: Comparison of sparsity degrees and smallest eigenvalues of UBG and PNP’s predictions with an Oracle GLasso, in both the strongly sparse (left panel) and weakly sparse (right) settings.
6 Conclusion & future directions
We have proposed Schur’s Positive-Definite Network (SpodNet), a novel learning module compatible with other standard architectures offering strict guarantees of SPD outputs. The principal novelty of SpodNet comes from its higher expressivity than existing approaches, as it is also able to handle additional desirable structural constraints. In this work, we highlighted that it can be used to efficiently learn sparse and SPD matrices. To the best of our knowledge, SpodNet layers are the first to offer strict guarantees of such highly non-trivial structure in the outputs. We have shown how to leverage these guarantees to build several neural architectures that outperform the standard solvers in the context of sparse precision matrix estimation. Current directions of research include algorithmic improvements in order to speed up SpodNet’s column/row/diagonal update framework and theoretical investigations in order to better understand the spectral perturbations undergone by the inputs.
7 Acknowledgments
We would like to thank Badr Moufad (Ecole Polytechnique, France) for his help with the PyTorch implementations of SpodNet, Paulo Gonçalves (Inria Lyon, France) for fruitful discussions as well as the Centre Blaise Pascal for computing ressources, which uses the SIDUS solution developed by Emmanuel Quemener (ENS Lyon, France) (Quemener and Corvellec, 2013). This work was partially supported by the AllegroAssai ANR-19-CHIA-0009 project.
Appendix A Supplementary Material
A.1 GLAD’s update rules
Each unrolled iteration of GLAD, which can be seen as an individual layer in a deep learning perspective, updates several running variables:
$\alpha^{+}=\tilde{f}(\|\bm{Z}-\bm{\Theta}\|_{F}^{2},\alpha)\,,$ (14)
$\bm{Y}^{+}=\frac{1}{\alpha^{+}}\bm{S}-\bm{Z}\,,$ (15)
$\bm{\Theta}^{+}=\frac{1}{2}\left(-\bm{Y}^{+}+\sqrt{{\bm{Y}^{+}}^{\top}\bm{Y}^{+}+\frac{4}{\alpha^{+}}\mathbf{Id}}\right)\,,$ (16)
$\lambda_{ij}^{+}=\tilde{h}(\Theta_{ij}^{+},S_{ij},Z_{ij})\,,$ (17)
$Z_{ij}^{+}=\operatorname{ST}_{\lambda_{ij}^{+}}\left(\Theta_{ij}^{+}\right)\,,\;\forall i,j\in[p]\,,$ (18)
in which the functions $\tilde{f}:\mathbb{R}^{2}\to\mathbb{R}$ and $\tilde{h}:\mathbb{R}^{3}\to\mathbb{R}$ are two small neural networks that are trained to predict adequate parameters for each layer. By construction, $\bm{\Theta}^{+}$ is always SPD and a sparsity structure is enforced on $\bm{Z}^{+}$, but the converse is not true.
A.2 Critical regularization parameter for the Graphical Lasso
Proposition A.1.
Let $\hat{\bm{\Theta}}=\operatorname*{argmin}_{\bm{\Theta}}-\log\det\bm{\Theta}+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{\Theta}\|_{1,\mathrm{off}}$. Then, $\hat{\bm{\Theta}}$ is diagonal if and only if $\lambda\geq\lambda_{\max}=\max_{i\neq j}|s_{ij}|$.
Proof.
The first-order optimality condition, aka Fermat’s rule, for the Graphical Lasso is:
$\hat{\bm{\Theta}}=\operatorname*{argmin}_{\bm{\Theta}}-\log\det\bm{\Theta}+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{\Theta}\|_{1,\mathrm{off}}\Leftrightarrow\hat{\bm{\Theta}}^{-1}+\bm{S}\in\partial\|\hat{\bm{\Theta}}\|_{1,\mathrm{off}}$ (19)
Now $\bm{U}\in\partial\|\hat{\bm{\Theta}}\|_{1,\mathrm{off}}$ if and only if
  • $U_{ii}=0$ for $i\in[p]$
  • $U_{ij}\in\lambda\partial|\Theta_{ij}|=\begin{cases}\lambda\operatorname{sign}(\Theta_{ij})$ if $\Theta_{ij}\neq 0\\ [-\lambda,\lambda]\text{ otherwise}\end{cases}$ for $i,j\in[p],i\neq j$.
Hence
[…]
$\Leftrightarrow s_{ij}\in[-\lambda,\lambda]\quad\forall i\neq j$ (21)
$\Leftrightarrow\lambda\geq\lambda_{\max}=\max_{i\neq j}|s_{ij}|$ (22)

### v3

@@
 While there exist various convex optimization-based estimators, they remain limited in expressivity due to their model-based approach.
-The success of deep learning has thus led many to use neural networks to learn to estimate SPD matrices in a data-driven fashion.
-For learning structured outputs, one promising strategy involves architectures designed by unrolling iterative algorithms, which potentially benefit from inductive bias properties.
-However, designing correct unrolled architectures for SPD learning is difficult: they either do not guarantee that their output has all the desired properties, rely on heavy computations, or are overly restrained to specific matrices which hinders their expressivity.
-In this paper, we propose a novel and generic learning module with guaranteed SPD outputs called SpodNet, that also enables learning a larger class of functions than existing approaches.
+The success of deep learning motivates the use of learning-based approaches to estimate SPD matrices with neural networks in a data-driven fashion.
+However, designing effective neural architectures for SPD learning is challenging, particularly when the task requires additional structural constraints, such as element-wise sparsity.
+Current approaches either do not ensure that the output meets all desired properties or lack expressivity.
+In this paper, we introduce SpodNet, a novel and generic learning module that guarantees SPD outputs and supports additional structural constraints.
 Notably, it solves the challenging task of learning jointly SPD and sparse matrices.
-Our experiments demonstrate the versatility of SpodNet layers.
+Our experiments illustrate the versatility and relevance of SpodNet layers for such applications.
 1 Introduction
-The estimation of symmetric positive-definite (SPD) matrices is a major area of research, due to their crucial role in various fields ranging such as optimal transport (Bonet et al., 2023; Ju and Guan, 2022; Yair et al., 2019), graph theory (Lauritzen, 1996), computer vision (Nguyen et al., 2019) or finance (Ledoit and Wolf, 2003).
+The estimation of symmetric positive-definite (SPD) matrices is a major area of research, due to their crucial role in various fields such as optimal transport (Bonet et al., 2023), graph theory (Lauritzen, 1996), computer vision (Nguyen et al., 2019) or finance (Ledoit and Wolf, 2003).
 While various statistical estimators, i.e.
-model-based, have been developed (Ledoit and Wolf, 2004; Banerjee et al., 2008; Cai et al., 2011), a more recent trend is to use learning-based approaches to estimate SPD matrices with neural networks (Gao et al., 2020).
+model-based, have been developed for specific tasks (Ledoit and Wolf, 2004; Banerjee et al., 2008; Cai et al., 2011), recent advancements focus on applying generic learning-based approaches to estimate appropriate SPD matrices with neural networks in a data-driven fashion (Huang and Van Gool, 2017; Gao et al., 2020).
 Training neural networks while enforcing non-trivial structural constraints such as positive-definiteness is a difficult task.
 There have been many efforts in this direction in recent years, often in an ad-hoc manner and each with their own shortcomings (see Section 2 for more details).
-Building on the seminal work of Gregor and LeCun (2010) in sparse coding, a promising research direction involves designing neural networks architectures from the unrolling of an optimization algorithm (Chen and Pock, 2016; Monga et al., 2021; Chen et al., 2022).
+Building on the seminal work of Gregor and LeCun (2010) in sparse coding, a promising research direction involves designing neural networks architectures from the unrolling of an optimization algorithm (Chen and Pock, 2016; Monga et al., 2021; Chen et al., 2022; Shlezinger et al., 2023).
 In the case of SPD matrices, algorithm unrolling presents several challenges.
@@
 These operations do not integrate well into a neural network architecture, making these algorithms difficult to unroll.
-Second, beyond ensuring positive definiteness, it is often desirable to impose additional structure on the matrix to estimate, such as elementwise sparsity.
-Simultaneously enforcing SPD + sparsity constraints on the output of a neural architecture is extremely difficult (Guillot and Rajaratnam, 2015; Sivalingam, 2015).
-So far, it is only a hypothetical property that is not satisfied in practice (see Figure 2 and Section 5 for more details).
-Contributions: In this paper, we combine the strengths of both model-based and data-driven approaches and we make the following contributions:
-  • We introduce a novel framework for building learning modules with a guaranteed SPD-to-SPD mapping for which it is also possible to enforce additional constraints on the output.
+Additionally, and more importantly, many applications require further structural constraints on the learned matrix.
+Elementwise sparsity is a typical example of such constraints (Banerjee et al., 2008), which significantly increases the complexity of the task: learning functions that simultaneously enforce SPDness and sparsity of the output is known to be challenging (Guillot and Rajaratnam, 2015; Sivalingam, 2015).
+There are currently no neural architectures that enable learning jointly SPD and sparse matrices.
+In this paper, we bridge this gap and make the following contributions:
+  • We introduce a new SPD-to-SPD neural network architecture that also supports enforcing additional constraints on the output.
 We refer to it as SpodNet for Schur’s Positive-Definite Network.
-Our approach is originally inspired by the block-matrix updates perspective of the Graphical Lasso solver (Mazumder and Hastie, 2012) and extended to generic SPD learning.
-As a particular case, SpodNet layers are able to learn jointly SPD and sparse matrices in a data-driven fashion.
-To the best of our knowledge, SpodNet-based architectures are the first to offer strict guarantees combining both properties.
-  • We illustrate the framework’s versatility through applications in sparse precision matrix estimation.
-In this scenario, we present limitations of other learning-based approaches and demonstrate how to overcome them with SpodNet layers.
-Our experiments validate the use of SpodNet for learning jointly SPD-to-SPD and sparsity inducing functions, and show promising results across various performance metrics.
+As a particular case, we show that SpodNet is able to learn jointly SPD and sparse matrices.
+To the best of our knowledge, SpodNet is the first architecture to provide strict guarantees for both properties.
+  • We demonstrate the framework’s relevance through applications in sparse precision matrix estimation.
+We highlight the limitations of other learning-based approaches and show how SpodNet addresses these issues.
+Our experiments validate SpodNet’s effectiveness in jointly learning SPD-to-SPD and sparsity-inducing functions, yielding competitive results across various performance metrics.
 Notation
@@
 On matrices, $\lVert\cdot\rVert_{1}$ is the sum of absolute values of the matrix coefficients.
-The cone of SPD matrices in dimension $p$ is denoted $\mathbb{S}^{p}_{++}$.
-We use $\bm{\Sigma}$ to designate covariance matrices, $\bm{\Theta}$ for precision matrices, and $\bm{S}$ for empirical covariance matrices.
+The cone of $p$ by $p$ SPD matrices is denoted $\mathbb{S}^{p}_{++}$.
+Figure 1: A SpodNet layer chains $p$ updates of column-row pairs and diagonals using neural networks. The matrices remain SPD at all times via Schur’s condition.
 2 Related works
-Before delving into the details of SpodNet, we briefly review existing approaches for constructing SPD matrices, which can be broadly divided into three categories.
-Riemannian learning
+We first review existing approaches for estimating SPD matrices, which can be broadly divided into three categories, all suffering the same limitation of not being able to handle additional structural constraints.
+Riemannian approaches
 Riemannian optimization provides tools to build algorithms whose iterates lie on Riemannian manifolds, such as the SPD manifold (Absil et al., 2008).
-Many have adopted these tools to design neural architectures that operate on the manifold of SPD matrices (Huang and Van Gool, 2017; Gao et al., 2020, 2022).
-However, a common and major bottleneck of these models is their dependence on Riemannian operations which are notoriously costly to compute.
-Furthermore, Riemannian optimization beyond smooth objective functions is still an active area of research (Absil and Hosseini, 2019; Huang and Wei, 2022; Zhou et al., 2023) and it is thus highly non-trivial to design neural architectures to handle additional constraints such as elementwise sparsity.
-Projection-based methods
-For some applications, it is of interest to project the matrices from a high-dimensional SPD manifold to a lower-dimensional one (Dong et al., 2017; Nguyen et al., 2019).
-Using linear algebra principles, these approaches design specific layers to guarantee that the final matrix is SPD.
-However, these solutions are ad-hoc and lack straightforward extensions for general SPD-to-SPD learning with additional structural constraints.
-Specifically, there is no obvious way to incorporate elementwise sparsity on the matrices without risking breaking their SPD structure.
-For instance, hard-thresholding the off-diagonal entries of a SPD matrix does not always preserve the SPD property (Guillot and Rajaratnam, 2012) and, more generally, elementwise functions preserving this property are very limited (Guillot and Rajaratnam, 2015).
-Unrolled learning
-In order to ensure that a network’s output strictly respects some desired properties, another recent and successful direction is to unroll convex optimization algorithms (Chen et al., 2022).
-This unrolling procedure acts as an “inductive bias” on the architecture (Shrivastava, 2020), naturally guiding the model to explore suitable solutions within the space of imposed properties.
+Many have adopted these tools to design neural architectures that operate on the manifold of SPD matrices (Huang and Van Gool, 2017; Gao et al., 2020; 2022).
+However, a common and significant bottleneck of these methods is the use of Riemannian operators, which are notoriously expensive to compute.
+Beyond this computational hindrance, and more importantly, current Riemannian methods are not able to impose additional sparsity on the learned matrices without breaking the SPD guarantee.
+SPD layers
+Neural layers with SPD outputs have also been proposed in Dong et al.
+(2017); Nguyen et al.
+(2019).
+A first line of work rely on the linear mapping $\bm{X}_{k+1}=\bm{W}_{k}\bm{X}_{k}\bm{W}_{k}^{\top}$ (with the layer’s weights $\bm{W}_{k}$ having full row-rank), while some others rely on clipping the eigenvalues of their output.
+Unfortunately, there are no obvious ways to incorporate additional structure such as elementwise sparsity on the matrices without risking breaking their SPD guarantee.
+For instance, hard-thresholding the off-diagonal entries of a SPD matrix does not in general preserve the SPD property (Guillot and Rajaratnam, 2012) and, more generally, elementwise functions preserving this property are very limited (Guillot and Rajaratnam, 2015).
+Unrolled neural architectures
+In order to ensure that a network’s output strictly respects some desired properties, a successful direction is to unroll convex optimization algorithms (Chen et al., 2022; Shlezinger et al., 2023).
+This unrolling procedure acts as an “inductive bias” on the architecture, naturally forcing the model to explore suitable solutions within the space of imposed properties.
 In SPD learning, this approach has been exploited by Shrivastava et al.
-(2019) who unrolled an optimization algorithm to train neural networks to estimate sparse inverses of covariance matrices.
-Whilst algorithm unrolling is a very powerful approach to learn specific matrices, it is very difficult to extend it to generic SPD learning.
-We provide more details about the limits of related work in this setting in Section 4.2.
+(2020) who unrolled an optimization algorithm to train neural networks to estimate inverses of covariance matrices.
+Whilst algorithm unrolling is a very powerful approach to learn specific matrices, it proves difficult to actually enforce several constraints simultaneously on these matrices.
+We provide further details about the limits of this approach in Section 4.1.
 3 The SpodNet framework
-We now introduce the SpodNet layer which is at the core of our contribution.
+We now introduce our core contribution, the SpodNet layer.
 Essentially, it is an SPD-to-SPD mapping parameterized by neural networks.
-As described in Section 4, SpodNet is inspired by an algorithm used to solve the Graphical Lasso in Mazumder and Hastie (2012).
-A SpodNet layer operates by cycling through the $p$ column-row pairs individually and i) updating the corresponding column-row with a neural network, and ii) updating the diagonal element to maintain the SPD constraint (see Figure 1).
-Interestingly, the neural network at step i) can update the column-row to any value without compromising the SPD guarantee.
-Consequently, one is free to exploit a spectrum of approaches for those updates depending on other desired structural properties of the output matrix.
-In particular, we describe in Section 4 a specific implementation of SpodNet that enforces elementwise sparsity for learning sparse precision matrices.
+A SpodNet layer operates by cycling through the $p$ column-row pairs individually by
+  1. (i)
+    updating the corresponding column-row with a neural network,
+  2. (ii)
+    updating the diagonal element to satisfy the SPD constraint (see Figure 1).
+Crucially, the neural network used at step (i) can update the column-row to any value without compromising the SPD guarantee.
+Consequently, one is free to exploit a spectrum of approaches for those updates, depending on other desired structural properties of the output matrix.
+In particular, we describe in Section 4 three specific implementations of SpodNet that enforce elementwise sparsity for learning sparse precision matrices.
 1:  Input: $\bm{\Theta}_{\mathrm{in}}\in\mathbb{S}_{++}^{p}$ and $\bm{W}_{\mathrm{in}}=\bm{\Theta}_{\mathrm{in}}^{-1}$
@@
 9:  Output: $\bm{\Theta}_{\mathrm{out}}\in\mathbb{S}_{++}^{p}$ and $\bm{W}_{\mathrm{out}}=\bm{\Theta}_{\mathrm{out}}^{-1}$
-Algorithm 1 One SpodNet layer
-Figure 1: Schematic overview of SpodNet layers.
-Our framework chains updates of column-row pairs and diagonals using neural networks (note: the updated columns and rows are not constant, despite being represented by constant colors).
-At all times the matrices remain SPD via Schur’s condition.
+Algorithm 1 The SpodNet layer
 3.1 Algorithmic foundations
-More formally, when the SpodNet layer operates on the $i$-th column and row of $\bm{\Theta}\in\mathbb{S}_{++}^{p}$ it also updates the $i$-th diagonal coefficient accordingly.
-The key to preserving the positive-definiteness of the matrix upon changing its $i$-th column and row is an appropriate update of the diagonal entry $\Theta_{ii}$ based on Schur’s condition for positive-definiteness.
+The key to preserving the positive-definiteness $\bm{\Theta}\in\mathbb{S}_{++}^{p}$ of the matrix upon after changing its $i$-th column and row is an appropriate update of the diagonal entry $\Theta_{ii}$ based on Schur’s condition for positive-definiteness.
 In the following, a $+$ superscript on a variable (e.g.
@@
 outputted by a neural network).
-The following proposition formalizes the SpodNet operations.
+The following proposition shows that SpodNet’s output is guaranteed to be SPD.
 Proposition 3.1.
@@
 $\bm{\Theta}=\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}_{12}\\ \bm{\theta}_{21}&\theta_{22}\end{bmatrix},\quad\text{with}\quad\bm{\Theta}_{11}\in\mathbb{R}^{(p-1)\times(p-1)},\,\bm{\theta}_{12}\in\mathbb{R}^{p-1},\ \bm{\theta}_{21}=\bm{\theta}_{12}^{\top},\,\theta_{22}\in\mathbb{R}\,.$ (1)
-(for a generic column $i$, $\bm{\Theta}_{11}$ refers to $\bm{\Theta}$ without its $i$-th row and $i$-th column, $\bm{\theta}_{12}$ is the $i$-th row of $\bm{\Theta}$ without its $i$-th value, and $\theta_{22}$ is $\Theta_{ii}$ as illustrated in Algorithm 1).
-Let $f:\mathbb{R}^{p\times p}\rightarrow\mathbb{R}^{p-1}$, $g:\mathbb{R}^{p\times p}\rightarrow\mathbb{R}_{++}$, where $f$ and $g$ are arbitrary functions.
+(for a generic column $i$, $\bm{\Theta}_{11}$ refers to $\bm{\Theta}$ without its $i$-th row and $i$-th column, $\bm{\theta}_{12}$ is the $i$-th row of $\bm{\Theta}$ without its $i$-th value, and $\theta_{22}$ is $\Theta_{ii}$ as illustrated below Algorithm 1).
+Suppose that $\bm{\Theta}\in\mathbb{S}_{++}^{p}$.
+Let $\bm{u}\in\mathbb{R}^{p-1}$ and $v>0$ be any vector and strictly positive scalar respectively.
 Then, updating the $i$-th row and column of $\bm{\Theta}$ as
-$\bm{\Theta}^{+}\triangleq\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}^{+}_{12}\triangleq f(\bm{\Theta})\\ \bm{\theta}_{21}^{+}\triangleq f(\bm{\Theta})^{\top}&\theta_{22}^{+}\triangleq g(\bm{\Theta})+\bm{\theta}^{+}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}^{+}_{12}\end{bmatrix}\,,$ (2)
+$\bm{\Theta}^{+}\triangleq\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}^{+}_{12}\triangleq\bm{u}\\ \bm{\theta}_{21}^{+}\triangleq\bm{u}^{\top}&\theta_{22}^{+}\triangleq v+\bm{u}^{\top}[{\bm{\Theta}}_{11}]^{-1}\bm{u}\end{bmatrix}\,,$ (2)
 preserves the SPD property, i.e. $\bm{\Theta}^{+}\in\mathbb{S}_{++}^{p}$.
+Since positivity of $\bm{\Theta}^{+}$ is guaranteed for any choice of $\bm{u}$ and $v$ as long as $v>0$, the principle of SpodNet updates is to learn these quantities as function of the current iterate, i.e.
+from the value of $\bm{\Theta}$, $\bm{\Theta}_{11}$, $\bm{\theta}_{12}$, etc (explicit forms for each variant will be given in Section 4.2).
+For simplicity, though they can depend on additional information, we write $\bm{u}=f(\bm{\Theta})$ and $v=g(\bm{\Theta})$, with $f$ and $g$ being learned mappings.
 Proof.
-Since $\bm{\Theta}$ is positive-definite, so is its leading principal submatrix $\bm{\Theta}_{11}$.
+Since $\bm{\Theta}$ is symmetric positive-definite, so is its leading principal submatrix $\bm{\Theta}_{11}$.
 It follows that $\bm{\Theta}^{+}$ is well-defined, and obviously symmetric.
-Its positive-definiteness then ensues from Schur’s condition for positive-definiteness.
-Indeed $\bm{\Theta^{+}}$ is SPD when $\bm{\Theta}_{11}\succ 0$ and $\theta_{22}^{+}-\bm{\theta}_{12}^{+\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}=g(\bm{\Theta})>0$ (Zhang, 2006, Theorem 1.12).
-The latter condition is ensured as the output of $g$ is a positive scalar.
+Its positive-definiteness ensues from Schur’s condition for positive-definiteness.
+Indeed $\bm{\Theta^{+}}$ is SPD when $\bm{\Theta}_{11}\succ 0$ and $\theta_{22}^{+}-\bm{\theta}_{12}^{+\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}>0$ (Zhang, 2006, Theorem 1.12).
+The latter condition is ensured as the left hand side is equal to $v$, which is strictly positive by assumption.
 ∎
-Finally, one SpodNet layer chains $p$ updates of the form (2), sequentially updating all column-row pairs one after the other as summarized in Algorithm 1.
-A full SpodNet architecture then stacks $K$ SpodNet layers.
-The overall procedure is schematized in Figure 1 for $K=2$.
-The expressivity of SpodNet comes from the simple realization that one is free to plug any arbitrary functions $f$ and $g$ in the updates whilst guaranteeing that $\bm{\Theta}$ remains in the SPD cone.
+Finally, one SpodNet layer chains $p$ updates of the form Equation 2, sequentially updating all column-row pairs one after the other as summarized in Algorithm 1.
+A full SpodNet architecture then stacks $K$ SpodNet layers, and, in practice, the first layer takes as input $\bm{\Theta}_{\mathrm{in}}=\left(\bm{S}+\mathbf{I}_{p}\right)^{-1}$ where $\bm{S}$ is the empirical covariance matrix.
+The overall architecture is schematized in Figure 1 for $K=2$.
+The expressivity of SpodNet comes from the flexibility to use any arbitrary functions $f$ and $g$ in the updates whilst guaranteeing that $\bm{\Theta}$ remains in the SPD cone.
 Namely, additional structure such as sparsity can trivially be imposed on $\bm{\Theta}$ through $f$.
 We next detail the computational cost of fully updating $\bm{\Theta}$ through one SpodNet layer.
-3.2 Improving update complexity
-The update of Equation 2 a priori requires inverting the matrix $\bm{\Theta}_{11}$ which comes with a prohibitive cost of $\mathcal{O}(p^{3})$ for each column-row update.
-We can fortunately leverage the column-row structure of the updates to decrease the cost to a mere $\mathcal{O}(p^{2})$, by maintaining the matrix $\bm{W}=\bm{\Theta}^{-1}$ up-to-date along the iterations.
+3.2 Improving the update complexity
+The diagonal update of Equation 2 a priori requires inverting the matrix $\bm{\Theta}_{11}$ which comes with a prohibitive cost of $\mathcal{O}(p^{3})$ for each column-row update.
+Fortunately, we are able to leverage the column-row structure of the updates to decrease the cost to a mere $\mathcal{O}(p^{2})$, by maintaining the matrix $\bm{W}=\bm{\Theta}^{-1}$ up-to-date along the iterations.
 Proposition 3.2.
@@
 Then $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$.
-In addition, if $\bm{\Theta}$ is updated as $\bm{\Theta}^{+}$ following Equation 2, then the update of $\bm{W}$ defined by
-$\bm{W}^{+}\triangleq\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}\bm{\theta}_{21}^{+}[{\bm{\Theta}}_{11}]^{-1}}{g(\bm{\Theta})}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}}{g(\bm{\Theta})}\\ \cdot&1/g(\bm{\Theta})\end{bmatrix}\,.$ (3)
+In addition, if $\bm{\Theta}$ is updated as $\bm{\Theta}^{+}$ following Equation 2 with $v=g(\bm{\Theta})$, then the update of $\bm{W}$ defined by
+$\bm{W}^{+}\triangleq\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}\bm{\theta}_{21}^{+}[{\bm{\Theta}}_{11}]^{-1}}{g(\bm{\Theta})}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}}{g(\bm{\Theta})}\\ \left(-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}^{+}}{g(\bm{\Theta})}\right)^{\top}&1/g(\bm{\Theta})\end{bmatrix}\,.$ (3)
 can be computed in $\mathcal{O}(p^{2})$ and satisfies $[\bm{W}^{+}]^{-1}=\bm{\Theta}^{+}$.
 Proof.
-By using the Banachiewicz inversion formula on Schur’s complement (Zhang, 2006, Theorem 1.2)
-$\begin{bmatrix}\bm{W}_{11}&\bm{w}_{12}\\ \bm{w}_{21}&w_{22}\end{bmatrix}$ $=\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}_{12}\\ \bm{\theta}_{21}&\theta_{22}\end{bmatrix}^{-1}=\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\\ \\ \cdot&\frac{1}{\theta_{22}-\bm{{\theta}}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\end{bmatrix}\,.$ (4)
+By the Banachiewicz inversion formula on Schur’s complement (Zhang, 2006, Thm 1.2),
+$\begin{bmatrix}\bm{W}_{11}&\bm{w}_{12}\\ \bm{w}_{21}&w_{22}\end{bmatrix}$ $=\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}_{12}\\ \bm{\theta}_{21}&\theta_{22}\end{bmatrix}^{-1}=\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\\ \\ \left(-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\right)^{\top}&\frac{1}{\theta_{22}-\bm{{\theta}}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\end{bmatrix}\,.$ (4)
 Identifying all blocks yields $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$ which can be computed in $\mathcal{O}(p^{2})$ if one has access to $\bm{W}$.
 This can be achieved using Equation 3, and involves only operations in $\mathcal{O}(p^{2})$.
-The property $\bm{\Theta}^{+}=[\bm{W}^{+}]^{-1}$ is statisfied by using the same inversion formula.
+The property $\bm{\Theta}^{+}=[\bm{W}^{+}]^{-1}$ is satisfied by using the same inversion formula.
 ∎
-Based on this result a full update of $\bm{\Theta}$ can be achieved with cost $\mathcal{O}(p^{3})$ ($p$ column-row updates of cost $\mathcal{O}(p^{2})$).
-4 Learning sparse precision matrices with SpodNet
-In this section, we propose three specific SpodNet-based architectures for learning sparse precision matrices, which naturally arise in the context of Gaussian graphical models (Lauritzen, 1996) that we now briefly recall.
-4.1 A model-based formulation: the Graphical Lasso
+A full update of $\bm{\Theta}$ can thus be achieved with cost $\mathcal{O}(p^{3})$ ($p$ column-row updates of cost $\mathcal{O}(p^{2})$).
+4 Using SpodNet to learn sparse precision matrices
+So far, we have not specified which choices of column and diagonal updates $f(\bm{\Theta})$ and $g(\bm{\Theta})$ we would use in practice.
+We now leverage the general framework of the SpodNet layer for learning SPD matrices with additional structure and propose three specific architectures for learning sparse and SPD matrices, with applications to sparse precision matrix learning.
+4.1 Inferring sparse precision matrices
 Consider a dataset of $n$ observed signals $\bm{x}_{1},\ldots,\bm{x}_{n}$ where each $\bm{x}_{i}\in\mathbb{R}^{p}$ follows a certain (centered) distribution with covariance matrix $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$.
 The problem of precision matrix estimation arises when attempting to identify the conditional dependency graph of the $p$ variables of this dataset.
-Estimating these dependencies is of interest in many fields such as neuroscience (Rubinov and Sporns, 2010), finance (Ledoit and Wolf, 2003) or genomics (Cai et al., 2013).
-In Gaussian graphical models, this graph is associated with the so-called precision matrix $\bm{\Theta}=\bm{\Sigma}^{-1}$ which directly encodes the dependencies [^1] In this case $\Theta_{ij}=0$ iff variable $i$ and $j$ are conditionally independent, given the other variables..
-The problem of precision matrix estimation involves determining $\bm{\Theta}\in\mathbb{S}^{p}_{++}$ given the empirical covariance matrix $\bm{S}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}$.
-Moreover, in many practical situations the graph (thus the matrix $\bm{\Theta}$) is also sparse due to limited conditional dependencies, which leads to a sparse precision matrix estimation problem.
-In this case, a very popular estimator is the Graphical Lasso (GLasso) (Banerjee et al., 2008; Friedman et al., 2008) that solves the convex optimization problem
+In Gaussian graphical models (Lauritzen, 1996), this graph is associated with the so-called precision matrix $\bm{\Theta}=\bm{\Sigma}^{-1}\in\mathbb{S}^{p}_{++}$.
+In this case $\Theta_{ij}=0$ iff the variables $i$ and $j$ are conditionally independent given the other variables.
+This estimation problem involves determining $\bm{\Theta}$ given the empirical covariance matrix $\bm{S}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}$.
+In most practical scenarios, the matrix $\bm{\Theta}$ is sparse due to limited conditional dependencies, which leads to a sparse SPD matrix estimation problem.
+For this problem, a very popular estimator is the Graphical Lasso (GLasso) (Banerjee et al., 2008; Friedman et al., 2008) that solves
 $\min_{\bm{\Theta}\succ 0}\ -\log\det(\bm{\Theta})+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{\Theta}\|_{1,\mathrm{off}}\,,$ (5)
 where $\|\bm{\Theta}\|_{1,\mathrm{off}}$ denote the off-diagonal $\ell_{1}$ norm of $\bm{\Theta}$, equal to $\sum_{i\neq j}|\Theta_{ij}|$.
-Equation 5 corresponds to an $\ell_{1}$-penalized maximum likelihood estimator under the assumption that the signals are Gaussian with covariance $\bm{\Sigma}$.
-It can be computed with various convex optimization solvers, e.g.
-primal formulations (Mazumder and Hastie, 2012), proximal gradient descent (Rolfs et al., 2012), or second order algorithms (Hsieh et al., 2014), all with strict guarantees of recovering both sparse and SPD precision matrices.
-In particular, the algorithm of Mazumder and Hastie (2012) precisely fits the form of the updates in (2), where the functions $f$ and $g$ are not learnt, but correspond to solutions of changing Lasso regression problems.
-Learning-based approaches, however, are scarcer.
-To the best of our knowledge, DeepGraph (Belilovsky et al., 2017) and GLAD (Shrivastava et al., 2019) are the only two existing works to tackle sparse precision matrix estimation under a data-driven perspective, both with restrictive limitations.
-This somewhat surprising fact is justified by the challenges posed by the task, which we highlight in the sequel.
-4.2 Limitations of learning-based approaches for sparse precision matrix estimation
-DeepGraph.Belilovsky et al.
-(2017) propose to directly map empirical covariance matrices to the graph structures using a CNN-based architecture trained with supervision.
-DeepGraph is trained for structure recovery only and not for the estimation of the actual precision matrix.
-It disregards the magnitude of the elements of the output matrix and focuses solely on its support, thus circumventing the problem of SPD learning, which is a limitation for some applications.
-Figure 2: GLAD limitations: Smallest eigenvalue (orange), density degree (green) and relative discrepancy of the two matrices (blue) for one of the output $(\bm{Z}_{\mathrm{out}},\bm{\Theta}_{\mathrm{out}})$ of GLAD.
-$\bm{\Theta}_{\mathrm{out}}$ and $\bm{Z}_{\mathrm{out}}$ are significantly different, $\bm{\Theta}_{\mathrm{out}}$ is not sparse, and $\bm{Z}_{\mathrm{out}}$ is not positive-definite.
-GLAD.
+It can be efficiently computed (Rolfs et al., 2012; Mazumder and Hastie, 2012; Hsieh et al., 2014) but suffers from the known limitations from the model-based approaches (Adler and Öktem, 2018; Arridge et al., 2019).
+To circumvent these limitations, two learning-based approaches have been proposed.
+Belilovsky et al.
+(2017) introduced DeepGraph, a CNN that directly maps empirical covariance matrices to the graph structures (i.e.
+to the support of the precision matrix).
+We emphasize that this model does not estimate $\bm{\Theta}$ but only its support.
 Shrivastava et al.
-(2019) unroll an alternating minimization algorithm for the Graphical Lasso (Dalal and Rajaratnam, 2017) to serve as an inductive bias for their model, which learns to predict elementwise regularization parameters of the updates.
-GLAD is trained with supervision, by minimizing a reconstruction error between the outputs and the ground truth matrix.
-The unrolled algorithm is designed to solve a Lagrangian relaxation of (5)
-$\min_{\bm{\Theta},\bm{Z}\in\mathbb{S}^{p}_{++}}-\log\det\left(\bm{\Theta}\right)+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{Z}\|_{1}+\frac{\alpha}{2}\|\bm{Z}-\bm{\Theta}\|_{F}^{2}\,.$ (6)
-As detailed in Section A.1, GLAD jointly operates on two matrices $\bm{Z}$ and $\bm{\Theta}$.
-GLAD is a powerful model, and has a minimalist yet effective parametrization.
-Unfortunately, by treating $\bm{Z}$ and $\bm{\Theta}$ separately, the architecture cannot learn matrices that are both sparse and SPD.
-This is illustrated in Figure 2 where $\bm{\Theta}$ and $\bm{Z}$ differ, $\bm{\Theta}$ is not sparse and $\bm{Z}$ is not SPD.
-4.3 New models: SpodNet for learning sparse precision matrices
-In this section, we present how to leverage SpodNet layers to overcome the limitations of existing solutions for learning both sparse and SPD precision matrices.
-By order of complexity, we present three neural architectures relying on SpodNet layers, each corresponding to a specific choice of the functions $f$ and $g$ in Equation 2.
-All three architectures are inspired by an unrolling of a proximal block coordinate descent applied to the Graphical Lasso problem (5), where the blocks are column-row pairs described in Section 3.
-Because (full) proximal gradient descent on (5) is called Graphical ISTA (Rolfs et al., 2012), we coin this Block-Graphical ISTA.
-To derive its iterations, we note that the gradient of $\bm{\Theta}\mapsto-\log\det(\bm{\Theta})+\langle\bm{S},\bm{\Theta}\rangle$ (the smooth term in the GLasso loss) is $-\bm{\Theta}^{-1}+\bm{S}$ (Boyd and Vandenberghe, 2004, Section A.4.1) and thus the gradient restricted to the $\bm{\theta}_{12}$-block is $-[\bm{\Theta}^{-1}]_{12}+\bm{s}_{12}=-\bm{w}_{12}+\bm{s}_{12}$.
-Consequently, the proximal gradient descent algorithm on these coordinates writes $\bm{\theta}_{12}^{+}=\mathrm{prox}_{\lambda\gamma\|\cdot\|_{1}}(\bm{\theta}_{12}-\gamma(\bm{s}_{12}-\bm{w}_{12}))$ where $\gamma>0$ is some stepsize and $\mathrm{prox}$ the proximal operator.
-Using that the soft-thresholding is the proximal operator of the $\ell_{1}$ norm, Block-Graphical ISTA hence reads
-$\bm{\theta}_{12}^{+}=\operatorname{ST}_{\gamma\lambda}(\bm{\theta}_{12}-\gamma(\bm{s}_{12}-\bm{w}_{12}))\,.$ (7)
+(2020) proposed GLAD, a neural architecture based on unrolling an alternating minimization algorithm for solving the GLasso.
+Based on a Lagrangian relaxation, GLAD iterates over two matrices $\bm{Z}$ and $\bm{\Theta}$ while learning step-size and thresholding parameters (see Section B.2 for more details).
+Out of these two matrices $\bm{Z}$ is sparse and $\bm{\Theta}$ is SPD, but GLAD fails to learn a single matrix that is both sparse and SPD.
+This limitation is highlighted in Figure 2 showing that, during training, $\bm{\Theta}$ and $\bm{Z}$ differ, $\bm{\Theta}$ is not sparse and $\bm{Z}$ is not SPD.
+Figure 2: GLAD’s limitations.
+Smallest eigenvalue (orange), density degree (green) and relative discrepancy of the two matrices (blue) for an output $(\bm{Z}_{\mathrm{out}},\bm{\Theta}_{\mathrm{out}})$ of GLAD.
+$\bm{\Theta}_{\mathrm{out}}$ and $\bm{Z}_{\mathrm{out}}$ are different, $\bm{\Theta}_{\mathrm{out}}$ is not sparse, and $\bm{Z}_{\mathrm{out}}$ is not positive-definite.
+4.2 SpodNet for sparse SPD learning
+We now show how SpodNet overcomes the limitations of current methods for learning both sparse and SPD matrices.
+By order of complexity, we introduce three new neural architectures, each corresponding to a specific choice of the functions $f,g$ for learning the values of $\bm{u}$ and $v$ in Equation 2.
+Our choices for $f,g$ are inspired by an unrolling of a proximal block coordinate descent applied to the GLasso problem, where the blocks are column-row pairs described in Section 3.
+A proximal coordinate gradient descent step on the GLasso objective updates the value of $\bm{\theta}_{12}$ with
+$\bm{\theta}_{12}^{+}=\operatorname{ST}_{\gamma\lambda}(\bm{\theta}_{12}-\gamma(\bm{s}_{12}-\bm{w}_{12}))\,,$ (6)
+where $\gamma>0$ is a step-size.
+This stems from the fact that the gradient of $\bm{\Theta}\mapsto-\log\det(\bm{\Theta})+\langle\bm{S},\bm{\Theta}\rangle$is $-\bm{\Theta}^{-1}+\bm{S}$ (Boyd and Vandenberghe, 2004, Section A.4.1), which restricted to the $\bm{\theta}_{12}$-block gives $-[\bm{\Theta}^{-1}]_{12}+\bm{s}_{12}=-\bm{w}_{12}+\bm{s}_{12}$.
+Together with the fact that the proximal operator of the $\ell_{1}$ norm is the soft-thresholding (Parikh et al., 2014) we get Equation 6.
+Because (full) proximal gradient descent on the GLasso is called Graphical ISTA (Rolfs et al., 2012), we coin Equation 6 Block-Graphical ISTA.
+We can now introduce our three architectures.
 Unrolled Block Graphical-ISTA (UBG)
-First, we propose to learn the stepsizes and soft-thresholding levels when unrolling the Block Graphical ISTA iterations (7). In a learning-based approach, we use as updating function $f$,
-$f_{\mathrm{UBG}}:\bm{\theta}_{12}\mapsto\operatorname{ST}_{\Lambda^{+}}(\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12}))\,,$ (8)
-where the elementwise soft-thresholding parameters $\Lambda^{+}$ along with the step-size $\gamma^{+}$ are predicted by perceptrons at each update.
-We emphasize that this is different from hyperparameter tuning, since $\gamma^{+}$ and $\Lambda^{+}$ are predicted at each update instead of being global parameters.
-UBG exhibits the highest inductive bias among the models, as it’s architecture is directly derived from unrolling an iterative optimization algorithm designed to minimize a model-based loss.
-Consequently, this architecture is particularly suited for learning sparse precision matrices.
-However, as highlighted in Section 3, the true power of SpodNet lies in the fact that one can plug any update functions and still get SPD outputs.
-This first motivates extending UBG to a more expressive setting with less inductive bias.
+First, we propose to learn the stepsizes and soft-thresholding levels when unrolling the Block Graphical ISTA iterations Equation 6. In a learning-based approach, we use as updating function $f$,
+$f_{\mathrm{UBG}}:\bm{\theta}_{12}\mapsto\operatorname{ST}_{\bm{\lambda}^{+}}(\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12}))\,,$ (7)
+where the step-size $\gamma^{+}=\operatorname{NN}_{1}(\bm{\theta}_{12})>0$ and the soft-thresholding parameters $\bm{\lambda}^{+}=\operatorname{NN}_{2}(\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12}))\in\mathbb{R}^{p-1}$.
+$\operatorname{NN}_{1}$ and $\operatorname{NN}_{2}$ are small multilayer perceptrons with architectures provided in Section A.2.
+We emphasize that this is different from hyperparameter tuning, since $\gamma^{+}$ and $\bm{\lambda}^{+}$ are predicted at each update instead of being global parameters.
+UBG exhibits the highest inductive bias among the models, as its architecture is directly derived from unrolling an iterative optimization algorithm designed to minimize a model-based loss.
+As highlighted in Section 3, one can plug any update functions and still get SPD outputs. This motivates extending UBG to a more expressive architecture.
 Plug-and-Play Block Graphical-ISTA (PNP)
-Next, we extend UBG to the Plug-and-Play setting (Venkatakrishnan et al., 2013; Romano et al., 2017; Kamilov et al., 2023).
-In a nutshell, these type of methods replace the proximal operator in first-order algorithms by a denoiser, usually implemented by a neural network.
-In our context, this corresponds to replacing the soft-thresholding of UBG by an arbitrary network $\Psi:\mathbb{R}^{p-1}\to\mathbb{R}^{p-1}$, that is,
+Precisely, we extend UBG to a Plug-and-Play-like setting (Venkatakrishnan et al., 2013; Romano et al., 2017).
+In a nutshell, these methods replace the proximal operator in first-order algorithms by a denoiser, usually implemented by a neural network.
+In our context, this corresponds to replacing the soft-thresholding of UBG by a neural network $\Psi:\mathbb{R}^{p-1}\to\mathbb{R}^{p-1}$, that is,
 $f_{\mathrm{PnP}}:\bm{\theta}_{12}\mapsto\Psi(\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12}))\,.$ (8)
 To promote sparsity of the output of $f_{\mathrm{PnP}}$, the last layer of $\Psi$ performs an elementwise soft-thresholding.
-The parameter for this soft-thresholding is also learned from the data and given by the same perceptron that predicts $\Lambda^{+}$ in UBG.
-End-to-end Learnable Block Graphical-ISTA (E2E)
+The parameter for this soft-thresholding is also learned from data and given by the same multilayer perceptron that predicts $\bm{\lambda}^{+}$ in UBG (Section A.2).
+End-to-end updates (E2E)
 Finally, we propose a fully-flexible architecture without any algorithm-inspired assumptions. Precisely, we consider
-$f_{\mathrm{E2E}}:\bm{\theta}_{12}^{+}\mapsto\Phi(\bm{\theta}_{12})\,,$ (10)
+$f_{\mathrm{E2E}}:\bm{\theta}_{12}\mapsto\Phi(\bm{\theta}_{12})\,,$ (9)
 where $\Phi$ takes in the current state of the column $\bm{\theta}_{12}$ and learns to predict an adequate column update $\bm{\theta}_{12}^{+}$.
 Intuitively, the neural network $\Phi$ acts as learning both the forward and the backward steps of a forward-backward iteration (Combettes and Pesquet, 2011).
-E2E is the most expressive of our three models because it is free from any inductive bias.
 As for PNP, sparsity of the predictions is enforced by a soft-thresholding non-linearity in the last layer of $\Phi$.
-Figure 3: Potential instabilities with no normalization: Along column update indices, we plot the smallest eigenvalue (blue), largest diagonal value (orange) and conditioning (green) of $\bm{\Theta}$.
-We observe that $\bm{\Theta}$ remains SPD but gets increasingly closer to being singular during the updates.
-Stability and update of the diagonal coefficient $\theta_{22}$
-Although the positive definiteness of $\bm{\Theta}$ is guaranteed to be preserved at each update for any positive-valued function $g$, we have empirically observed that its smallest eigenvalue could approach 00 as visible in Figure 3.
+Figure 3: Training dynamics of our 3 models (on test data described in Section 5.1).
+Left: The outputs of our 3 models remain positive-definite.
+Middle: The conditioning remains stable.
+Right: The outputs are sparse.
+Overall our models produce jointly sparse + SPD outputs.
+Finally, for all models, we use for $g$ a small neural network that takes as input $\theta_{22},s_{22}$ and $(\bm{\theta}_{12}^{+})^{\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}$.
+Its positivity is ensured by using an absolute value function as final nonlinearity.
+Improving training stability
+Although the positive definiteness of $\bm{\Theta}$ is guaranteed to be preserved at each update for any positive-valued function $g$, we have empirically observed that its smallest eigenvalue could approach 00 as visible in Section B.1.
 In practice this leads to instability in training.
-Below, we offer an interpretation of this phenomenon and suggest a solution to address it.
+Below, we provide an interpretation of this phenomenon and provide a solution to address it.
 Each column-row update inside a SpodNet layer can be written as the rank-$2$ update
@@
 Experimentally, the most successful approach to control $\|\Delta_{\bm{\Theta}}\|_{\mathrm{op}}$ is to limit the magnitude of $\theta^{+}_{22}$, which is not a trivial task as the updated value $\theta^{+}_{22}=g(\bm{\Theta})+\bm{\theta}^{+}_{21}[\bm{\Theta}_{11}]^{-1}\bm{\theta}^{+}_{12}$ must satisfy $\theta^{+}_{22}-\bm{\theta}^{+}_{21}[\bm{\Theta}_{11}]^{-1}\bm{\theta}^{+}_{12}>0$.
-Thus, in practice we propose to scale $\bm{\theta}_{12}^{+}$, ultimately choosing for all 3 models $f\in\{f_{\mathrm{UBG}},f_{\mathrm{PnP}},f_{\mathrm{E2E}}\}$ the column update
-$\bm{\theta}_{12}^{+}=\sqrt{\zeta}\frac{f(\bm{\Theta})}{\sqrt{f(\bm{\Theta})^{\top}[\bm{\Theta}_{11}]^{-1}f(\bm{\Theta})}}\,,$ (13)
-which ensures that the diagonal update is given by $\theta_{22}^{+}=g(\bm{\Theta})+\zeta$.
-The hyperparameter $\zeta>0$ acts as a form of regularization that handles a compromise between stability and expressivity of the model.
-In all of our experiments we set $\zeta=1$.
-Finally, for all models, we use for $g$ a small neural network that takes as input the previous value of $\theta_{22}$ and the corresponding diagonal entry of $\bm{S}$, $s_{22}$, and whose positivity is ensured by using an exponential function as final nonlinearity.
+Thus, we propose to scale $\bm{\theta}_{12}^{+}$ by scaling the preactivation [^1] Experimentally scaling the preactivation works better than scaling the network’s output.
+$\bm{z}$ of the last layer of $f$ by $\frac{\sqrt{\zeta}}{\sqrt{\bm{z}^{\top}[\bm{\Theta}_{11}]^{-1}\bm{z}}}$.
+The hyperparameter $\zeta>0$ acts as a form of regularization that handles a compromise between stability and expressivity of the model; in all of our experiments we use $\zeta=1$.
+As visible in Figure 3 this scaling ensures a smooth training of the network, with stable condition number of the $\bm{\Theta}$ matrix.
 5 Experiments
-We now illustrate our three graph learning models built on SpodNet layers’s ability to learn relevant sparse precision matrices.
-This section aims to achieve two objectives.
-First, we highlight that our three models learn matrices that are indeed sparse and SPD in contrast to other learning-based approaches such as GLAD.
-Second, we show that our models identify accurate sparse precision matrices and even outperform an oracle Graphical Lasso estimator.
-Figure 4: Training dynamics of our UBG, PNP and E2E models, compared to the baseline Oracle GLasso (in red) which provides both sparse and SPD matrices.
-Data
-We use synthetically generated data to fit our models.
-We generate $N$ sparse SPD matrices of dimension $p=50$ using sklearn’s make_sparse_spd_matrix function (Pedregosa et al., 2011), of which we ensure proper conditioning by adding $0.1$ to the diagonal entries.
+We now illustrate SpodNet’s ability to learn SPD matrices with additional structure by using our three derived graph learning models (UBG, PNP and E2E) to learn sparse precision matrices, on both synthetic and real data.
+Additional details regarding the experimental setups can be found in Appendix A.
+PyTorch implementations are provided with the submission and will be made public upon acceptance.
+5.1 Sparse precision matrix recovery
+Our goal in this section is to evaluate our models’ performance and generalization ability regarding the recovery of sparse precision matrices on synthetic data.
+The following experiments serve several purposes: (1) compare learning-based methods against traditional model-based methods, (2) evaluate our models’ reconstruction performance in terms of matrix estimation and support recovery.
+Data & training
+We use synthetic data to train our SpodNet models as in Belilovsky et al.
+(2017); Shrivastava et al.
+(2020).
+We generate $N$ sparse SPD $p\times p$ matrices using sklearn’s make_sparse_spd_matrix function (Pedregosa et al., 2011), of which we ensure proper conditioning by adding $0.1\cdot\mathbf{I}_{p}$.
 These matrices are treated as ground truth precision matrices $\bm{\Theta}_{\mathrm{true}}^{(i)}$ for $i\in[N]$.
-The sparsity degree of each matrix is controlled through the one imposed on their Cholesky factors during their generation.
-We consider two setups for these sparse SPD matrices: weakly sparse ones ($\alpha=0.7$), and strongly sparse ones ($\alpha=0.95$).
-From each of these $N$ ground truth precision matrices $\bm{\Theta}_{\mathrm{true}}=\bm{\Sigma}_{\mathrm{true}}^{-1}$, we sample $n=50$ i.i.d.
-centered Gaussian random vectors $\bm{x}_{j}\sim\mathcal{N}(0,\bm{\Sigma}^{(i)}_{\mathrm{true}}),j\in[n]$, which are used to compute an empirical covariance matrix $\bm{S}^{(i)}$.
-Both the training and testing datasets thus comprise distinct couples $(\bm{S}^{(i)},\bm{\Theta}^{(i)}_{\mathrm{true}})$, meaning that each $\bm{S}^{(i)}$ stems from a different distribution.
-Since $n=p$, the $\bm{S}^{(i)}$’s are poorly conditioned and the estimation of $\bm{\Theta}_{\mathrm{true}}^{(i)}$ is typically considered to be a very difficult problem to solve.
-We generate $N_{\mathrm{test}}=100$ $(\bm{\Theta},\bm{S})$ couples for the testing set on which we evaluate all approaches.
-The average sparsity degrees of the ground truth in the weakly sparse and strongly sparse testing sets are respectively about 24% and 90%.
-The training set for our UBG model comprises 100 couples, whereas PNP and E2E are trained on 1000 couples.
-Minimization objective & performance metric
-All three models are trained using ADAM with default parameters (Kingma and Ba, 2014) to minimize a reconstruction error loss $\mathcal{L}_{\mathrm{MSE}}=\frac{1}{N_{\mathrm{train}}}\sum^{N_{\mathrm{train}}}_{i=1}\|\hat{\bm{\Theta}}^{(i)}-\bm{\Theta}^{(i)}_{\mathrm{true}}\|^{2}_{F}$ in the vein of GLAD, where $N_{\mathrm{train}}$ is the size of the training set and $\hat{\bm{\Theta}}$ is the model’s output.
-Architectures & training
-We consider in our experiments models with $K=1$ SpodNet layer.
-The input to the models are the empirical covariance matrices $\bm{S}$, from which we compute $\bm{\Theta}_{\mathrm{init}}=(\bm{S}+0.1*\bm{I})^{-1}$ and $\bm{W}_{\mathrm{init}}=\bm{S}+0.1*\bm{I}$, which serve as the inputs to the SpodNet layer.
-We use learning rates of $0.01$, $0.02$ and $0.03$ respectively, to train UBG, PNP and E2E.
-The principal difficulty in training UBG and PNP stems from learning to predict $\gamma^{+}$.
-We find that using a perceptron with a single hidden layer of 10 neurons and ReLU activations that takes in the current state of the whole matrix to predict a $\gamma^{+}$ is expressive enough to learn adequate matrices (Figures 4 and 6).
-Its positivity is ensured by taking the absolute value of the prediction.
-We believe that a more extensive architectural search for this perceptron could help achieve even better results.
-The prediction of the $\Lambda^{+}$ parameters at each step is performed by a small single layer perceptron of 5 neurons with ReLU activation functions and an absolute value function following the last layer to ensure positivity.
-This network acts elementwise on the entries of the input vector.
-The sufficiency of a lower amount of training data for UBG could be imputed to its stronger inductive bias.
-PNP’s $\Psi$ function and E2E’s $\Phi$ function are learned by a single layer ReLU perceptron of 500 neurons.
-All experiments were ran on an RTX 2080 SUPER GPU with an AMD EPYC 7302 CPU and resources of similar computing power.
-Sparsity & positive-definiteness
-Figure 5 highlights that all three of our models are indeed able to learn jointly sparse and SPD matrices.
-Contrary to GLAD, the raw predictions respect these structural constraints by construction.
-For the sake of completeness, we threshold the entries of GLAD’s raw predictions at increasing values to artificially introduce sparsity, and observe that this can indeed break their positive-definiteness.
-As visual confirmation, Figure 6 shows one of UBG and PNP’s predictions in the weakly sparse and strongly sparse settings.
-Performance & benchmark
-Since GLAD’s predictions are not naturally sparse and SPD, we benchmark against sklearn’s implementation of the GLasso solver (Friedman et al., 2008), which is the standard algorithm for learning sparse precision matrices.
-We compare ourselves to an ideal Oracle GLasso where the $\ell_{1}$ hyperparameter $\lambda^{(i)}$ is individually assigned to the solver for each target matrix $\bm{\Theta}_{\mathrm{true}}^{(i)}\forall i\in[N_{\mathrm{test}}]$ and where $\lambda^{(i)}$ is optimally tuned so as to achieve best possible reconstruction error.
-In practice, we validate this $\lambda^{(i)}$ in $[10^{-3}\lambda_{\max},\lambda_{\max}]$ where $\lambda_{\max}$ is the parameter such that every off-diagonal entry of the prediction is zero (see Section A.2).
-We emphasize that this is an idealistic scenario where the GLasso is at its most favorable.
-The results are depicted in Figure 4 for the weakly sparse setting and illustrates that our three models are indeed able to learn jointly sparse and SPD matrices while outperforming the Oracle GLasso solver.
-PNP yields the best performance in terms of NMSE whereas the E2E model is best in terms of support recovery in our setting.
-Figure 6 offers visual confirmation of the relevancy of the predictions.
-Interestingly, merely only one SpodNet layer is enough to achieve good performances for all of the three models.
-Figure 5: Sparsity (left) and positive-definiteness (right) of the raw predictions of all three of our models and GLAD’s, along with the latter’s thresholded predictions, on the testing set of the weakly sparse setting.
-Figure 6: Comparison of sparsity degrees and smallest eigenvalues of UBG and PNP’s predictions with an Oracle GLasso, in both the strongly sparse (left panel) and weakly sparse (right) settings.
-6 Conclusion & future directions
-We have proposed Schur’s Positive-Definite Network (SpodNet), a novel learning module compatible with other standard architectures offering strict guarantees of SPD outputs.
-The principal novelty of SpodNet comes from its higher expressivity than existing approaches, as it is also able to handle additional desirable structural constraints.
-In this work, we highlighted that it can be used to efficiently learn sparse and SPD matrices.
+The sparsity degree of each matrix is controlled through the one imposed on their Cholesky factors during their generation: in the strongly sparse setting $\bm{\Theta}_{\mathrm{true}}$ has roughly 90 % of zero entries while in the weakly sparse setting this value is around 25 %.
+For each of these $N$ ground truth precision matrices $\bm{\Theta}_{\mathrm{true}}$, we sample $n$ i.i.d.
+centered Gaussian random vectors $\bm{x}_{j}\sim\mathcal{N}(0,(\bm{\Theta}^{(i)}_{\mathrm{true}})^{-1})$, which are used to compute an empirical covariance matrix $\bm{S}^{(i)}$.
+The dataset thus comprises couples $(\bm{S}^{(i)},\bm{\Theta}^{(i)}_{\mathrm{true}})$ where each $\bm{S}^{(i)}$ stems from a different ground truth precision matrix.
+For all the experiments, we generate $N_{\mathrm{train}}=1000$ different $(\bm{S},\bm{\Theta}_{\mathrm{true}})$ couples for the training set and $N_{\mathrm{test}}=100$ couples for the testing set on which we validate our models.
+Further details on the parameters used during the data generation are in Section A.1.
+We train our three models (UBG, PNP and E2E) to minimize a reconstruction error $\mathcal{L}_{\mathrm{MSE}}=\frac{1}{N_{\mathrm{train}}}\sum^{N_{\mathrm{train}}}_{i=1}\|\hat{\bm{\Theta}}^{(i)}-\bm{\Theta}^{(i)}_{\mathrm{true}}\|^{2}_{F}$ in the vein of Gregor and LeCun (2010); Shrivastava et al.
+(2020), where $\hat{\bm{\Theta}}$ is the model’s output.
+All three models are trained using ADAM with default hyperparameters (Kingma and Ba, 2014).
+Baselines
+We compare our three models to various other approaches throughout our experiments: learning-based, Riemannian and model-based estimators.
+As performed in Belilovsky et al.
+(2017); Shrivastava et al.
+(2020), we use the three most popular methods for estimating precision matrices as our core baselines: the GLasso (Friedman et al., 2008), the inverse Ledoit-Wolf estimator (Ledoit and Wolf, 2004) and the inverse OAS (Chen et al., 2010).
+The Ledoit-Wolf and OAS estimators are covariance matrix estimators, for which we take the inverse for estimating the precision (we simply refer to them as Ledoit-Wolf and OAS).
+All three approaches guarantee the positive-definiteness of the estimated matrix but the GLasso is the only one to additionally ensure sparse predictions.
+For the GLasso, for each matrix $\bm{S}^{(i)}$ a sparsity-regulating hyperparameter $\lambda^{(i)}$ is chosen using sklearn’s GraphicalLassoCV cross-validation procedure, that relies only on the samples used in $\bm{S}^{(i)}$.
+In the same vein, we use dedicated Ledoit-Wolf and OAS estimators for each $\bm{S}^{(i)}$.
+We also compare our approaches to the GLAD model (Shrivastava et al., 2020), considering both its $\bm{\Theta}$ and $\bm{Z}$ outputs.
+We recall that $\bm{Z}$ is not guaranteed to be positive-definite and its validity as a precision matrix estimator is therefore debatable.
+GLAD is trained with the same loss, data and optimizer as our models.
+We compare the performance of the different approaches in terms of Normalized MSE on the test set: $\mathrm{NMSE}=\frac{1}{N_{\mathrm{test}}}\sum^{N_{\mathrm{test}}}_{i=1}\frac{\|\hat{\bm{\Theta}}^{(i)}-\bm{\Theta}^{(i)}_{\mathrm{true}}\|^{2}_{F}}{\|\bm{\Theta}^{(i)}_{\mathrm{true}}\|^{2}_{F}}$ and F1 score for assessing support recovery.
+For the latter, false positives correspond to $\hat{\bm{\Theta}}_{ij}\neq 0$, $(\bm{\Theta}_{\mathrm{true}})_{ij}=0$ (and similarly for true positives and false negatives).
+As shown in Figure 4, at convergence, the learning-based methods show to be more suited for the actual reconstruction of $\bm{\Theta}_{\mathrm{true}}$ as they significantly outperform traditional methods.
+This superior performance is especially pronounced as the dimensionality of the data grows and the sparsity of the ground truth decreases.
+In the most challenging scenario ($p=100$ and $n=100$ with weakly sparse $\bm{\Theta}_{\mathrm{true}}$), the traditional methods perform poorly, with NMSE around 90%.
+Figure 4: Learning-based (in variations of blue) vs traditional methods (in variations of red).
+Dotted curves indicate when one of the constraints (SPDness or sparsity) is not guaranteed.
+First row: Strongly sparse $\bm{\Theta}_{\mathrm{true}}$.
+Second row: Weakly sparse $\bm{\Theta}_{\mathrm{true}}$.
+Figure 5: Comparing models in sample deficient regimes ($n\leq p$) up to large-sample regimes ($n\gg p$), evaluated in terms of NMSE and F1 score for support recovery.
+Dotted curves indicate when one of the constraints (SPDness or sparsity) is not guaranteed.
+In large dimension, GLAD’s $\bm{Z}$ performs well, but is never SPD; GLAD’s $\bm{\Theta}$ is SPD, but performs badly.
+Figure 4 also suggests that our models perform especially well in terms of NMSE in the low-samples regime ($n\leq p$), which is a common setting in various real-world applications, such as neuroscience time-series analysis or financial temporal network inference.
+In Figure 5, we thus study the influence of the number of samples $n$ on the performance of each method.
+The results show that although all methods achieve excellent recovery in the large sample regime (down to around 2.5% NMSE in the lower dimensional $p=20$ setting), our models are especially more suited for matrix reconstruction in case of sample-deficiency, especially so in the higher dimensional ($p=100$) setting.
+In terms of support recovery, our models achieve significantly better F1 scores than the GLasso as the number of sample $n$ grows, with up to 25% improvement in certain tested settings.
+In the lower dimensional setting ($p=20$), we are even able to achieve a F1 score of over 0.75 with our UBG model.
+Finally, although GLAD’s $\bm{Z}$ performs well, an additional experiment in Section B.3 shows that those matrices turn out to never be SPD in various settings, making them unsuited as precision matrix estimators.
+5.2 Application to unsupervised graph learning on a real-world dataset
+We finally evaluate the performance of SpodNet in a graph learning context using a real-world dataset and under an unsupervised scenario.
+The aim of this experiment is twofold: (1) to assess the generalization capabilities of SpodNet, and (2) to determine if it can yield an accurate graph topology when trained solely on synthetic data, following the idea of Belilovsky et al.
+(2017).
+Figure 6: Graphs estimation on the Animals dataset with the GLasso, EGFM and our UBG model.
+The former and the latter yield cleaner and more interpretable clusters than the GLasso, but EGFM is specifically tailored for this task as opposed to our UBG model.
+Data & training
+We consider the Animals dataset (Lake and Tenenbaum, 2010), which comprises $p=33$ animal species, each characterized by responses to $n=102$ binary questions (e.g., "Has teeth?", "Is poisonous?").
+The objective is to infer a graph of connections between these $p$ animals based on the $n$ responses.
+We train our UBG model by minimizing the MSE on the synthetic data detailed in Section 4.2, with $p=33,n=102$ and various levels of sparsity.
+We generate 1000 training matrices and train the model for 100 epochs.
+The predicted precision matrix is then computed by inputting the empirical covariance matrix from the Animals dataset into SpodNet.
+Baselines
+We compare our approach with the GLasso (with the regularization parameter selected by cross-validation), and the Elliptical Graphical Factor Model (EGFM) proposed by Hippert-Ferrer et al.
+(2023).
+The latter is a Riemannian optimization method specifically designed to identify clusters within the data by estimating a precision matrix whose inverse is modeled as a low-rank matrix plus a positive diagonal.
+For a fair comparison, we adopt the same hyperparameters as outlined in the original paper, setting the rank and regularization parameter to 10.
+Each method yields a precision matrix, which is used to construct a graph representing the connections between the $p$ animals.
+The graph is constructed by considering the absolute values of the precision matrix coefficients, which represent the strength of the connections, and by removing the diagonal elements to eliminate self-loops.
+The results are illustrated in Figure 6, where the nodes of the graphs are partitioned using the Louvain algorithm (Blondel et al., 2008).
+Qualitatively, we observe that UBG produces a coherent graph, with outcomes comparable to those of GLasso and EGFM.
+Additionally, the graph learned by our UBG model exhibits a cleaner structure compared to GLasso, with a sparser representation and finer clusters, such as the grouping of (Chimp, Gorilla).
+To quantitatively assess the quality of the obtained graphs, we calculate the modularity $m$ of the partitions: higher modularity values indicate better separation of the graph into distinct subcomponents (Newman, 2006).
+We find that the graph produced by UBG achieves a higher modularity ($m=0.78$) than GLasso ($m=0.61$) and slightly lower than EGFM ($m=0.86$).
+It is important to note that the EGFM is specifically designed to obtain clustered graphs, whereas our UBG model has no such inductive bias.
+These combined quantitative and qualitative results demonstrate that SpodNet generalizes effectively to unseen data and produces a coherent graph structure.
+6 Conclusion
+We have proposed Schur’s Positive-Definite Network (SpodNet), a novel learning module compatible with other standard architectures, offering strict guarantees of SPD outputs.
+The principal novelty of SpodNet comes from its ability to handle additional desirable structural constraints, such as elementwise sparsity which we used as an illutrative example throughout this paper.
 To the best of our knowledge, SpodNet layers are the first to offer strict guarantees of such highly non-trivial structure in the outputs.
-We have shown how to leverage these guarantees to build several neural architectures that outperform the standard solvers in the context of sparse precision matrix estimation.
-Current directions of research include algorithmic improvements in order to speed up SpodNet’s column/row/diagonal update framework and theoretical investigations in order to better understand the spectral perturbations undergone by the inputs.
+In future works, SpodNet could be leveraged to learn other additional structures beyond sparsity.
+We have shown how to leverage SpodNet to build neural architectures that outperform traditional methods in the context of sparse precision matrix estimation in various settings.
+Future directions of research include theoretical investigations in order to better understand the expressivity and derive convergence guarantees for our approach.
+7 Reproducibility statement
+We believe this paper fully discloses the information needed to reproduce the main experimental results.
+Our core contribution is the SpodNet layer: its general algorithmic framework is presented in Algorithm 1.
+Three specific models built using this framework for learning sparse and SPD matrices are introduced in Section 4.2, and the exact architectures used for each of them throughout our experiments are explained in detail in Section A.2.
+The conducted experiments are described in Section 5 with additional details in Appendix A.
+PyTorch implementations are provided along with the submission.
 8 Acknowledgments
-We would like to thank Badr Moufad (Ecole Polytechnique, France) for his help with the PyTorch implementations of SpodNet, Paulo Gonçalves (Inria Lyon, France) for fruitful discussions as well as the Centre Blaise Pascal for computing ressources, which uses the SIDUS solution developed by Emmanuel Quemener (ENS Lyon, France) (Quemener and Corvellec, 2013).
+We would like to thank Badr Moufad (Ecole Polytechnique, France) for his help with the implementations, Paulo Gonçalves (Inria Lyon, France) for fruitful discussions as well as the Centre Blaise Pascal for computing ressources, which uses the SIDUS solution developed by Emmanuel Quemener (ENS Lyon, France) (Quemener and Corvellec, 2013).
 This work was partially supported by the AllegroAssai ANR-19-CHIA-0009 project.
-Appendix A Supplementary Material
-A.1 GLAD’s update rules
-Each unrolled iteration of GLAD, which can be seen as an individual layer in a deep learning perspective, updates several running variables:
+Appendix A Experimental details
+A.1 Synthetic data generation
+Here we provide additional details on the experiment of Section 5.1.
+The sparsity degree of each matrix is controlled through the one imposed on their Cholesky factors during their generation, with the alpha parameter of the function.
+Throughout our experiments, we consider alpha $\in\{0.7,0.95\}$; this is the probability of a 00 entry on the Cholesky factor of the generated matrix and not the actual fraction of null elements of $\bm{\Theta}$.
+Although the numbers vary with the dimension $p$, $\alpha=0.95$ roughly leads to 90 % of zero entries, while $\alpha=0.7$ leads to 25 % of zero entries.
+A.2 Architectures for UBG, PNP and E2E
+The function $g$ for all three models is a MLP with two hidden layers of 3 neurons each with a ReLU activation function.
+It takes as input $\theta_{22},s_{22}$ and $(\bm{\theta}_{12}^{+})^{\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}$.
+UBG
+Its $\gamma^{+}$ parameter is predicted by a MLP that takes as input the current state of the $\bm{\theta}_{12}$ block (of dimension $p-1$) and has a single hidden layer of $\lfloor p/2\rfloor$ neurons with a ReLU activation function.
+The (single) output neuron has an absolute value activation, in order to keep it positive since it predicts a step-size.
+UBG involves another MLP to learn the vector $\bm{\lambda}^{+}$ of elementwise soft-thresholding parameters.
+This MLP has a single hidden layer of 5 neurons, with ReLU activation function.
+For the same reason as before, the output layer also has an absolute value activation.
+PNP
+First, the step-size $\gamma^{+}$ is predicted by a MLP with the exact same architecture as the one in UBG.
+The learned operator $\Psi$ takes as input $\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12})$, passes it through a single hidden layer of $2p$ neurons, followed by a ReLU activation function, and projects it back into a $p-1$ dimensional vector.
+The architecture to predict $\bm{\lambda}$ is the same as for UBG. Our experiments show that multiplying its output by $0.1$ helps avoid local minimas and improves convergence.
+E2E
+The neural network $\Phi$ is a MLP that takes as input $\bm{\theta}_{12}$, passes it through a single hidden layer of $10p$ neurons follow by a ReLU activation function, and projects it back into a $p-1$ dimensional vector.
+The architecture to predict $\bm{\lambda}$ is the same as for UBG and PNP.
+Our experiments show that multiplying its output by $0.1$ helps avoiding local minimas and improves convergence.
+GLAD
+We use the authors’ original implementation retrieved from the authors repository at https://github.com/Harshs27/GLAD, with the default hyperparameters.
+All four models are implemented with $K=1$ layer throughout all the experiments, since our results show that this was enough to yield competitive to outperforming results in the settings under consideration.
+A.3 Details on Figures 4 and 5
+For the weakly sparse settings, we use a learning rate of $10^{-3}$ for all three of our own models.
+For the strongly sparse settings, we use a learning rate of $10^{-2}$.
+GLAD’s learning rate is set to $10^{-2}$ in all settings.
+All four models are trained on the same 1000 training matrices, using a batch-size of 10, with ADAM’s default parameters.
+The models are tested on the same 100 matrices as mentioned in Section 5.1.
+For Figure 5 We use the strongly sparse ground truths in order for the F1 score to be more sensical, and to better illustrate the relevancy of the learned sparsity patterns.
+We use a learning rate of $10^{-2}$ for all three of them in every setting for $p=100$, and of $5\cdot 10^{-3},10^{-2},10^{-3},3\cdot 10^{-2}$ in the $p=20$ setting for respectively $n=10,n=20,n=100,n=200$ and $n=500$.
+Appendix B Additional details on models
+B.1 SpodNet’s potential instabilities without normalization
+We show in Figure 7 the evolution of the conditioning of an example of $\bm{\Theta}$ during training that becomes unstable if we do not use the stabilization procedure described in Section 4.
+We observe that although the matrix remains SPD as predicted by Proposition 3.1 but gets increasingly closer to being singular during the updates.
+Figure 7: Potential instabilities without normalization.
+Along column update indices, we plot the smallest eigenvalue (blue), largest diagonal value (orange) and conditioning (green) of a $\bm{\Theta}$.
+B.2 GLAD’s update rules
+GLAD unrolls an algorithm that solves the following optimization problem:
+$\min_{\bm{\Theta},\bm{Z}\in\mathbb{S}^{p}_{++}}-\log\det\left(\bm{\Theta}\right)+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{Z}\|_{1}+\frac{\alpha}{2}\|\bm{Z}-\bm{\Theta}\|_{F}^{2}\,.$ (12)
+Each iteration of GLAD, which can be seen as an individual layer in a deep learning perspective, updates several running variables:
 $\alpha^{+}=\tilde{f}(\|\bm{Z}-\bm{\Theta}\|_{F}^{2},\alpha)\,,$ (13)
@@
 By construction, $\bm{\Theta}^{+}$ is always SPD and a sparsity structure is enforced on $\bm{Z}^{+}$, but the converse is not true.
-A.2 Critical regularization parameter for the Graphical Lasso
-Proposition A.1.
-Let $\hat{\bm{\Theta}}=\operatorname*{argmin}_{\bm{\Theta}}-\log\det\bm{\Theta}+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{\Theta}\|_{1,\mathrm{off}}$.
-Then, $\hat{\bm{\Theta}}$ is diagonal if and only if $\lambda\geq\lambda_{\max}=\max_{i\neq j}|s_{ij}|$.
-Proof.
-The first-order optimality condition, aka Fermat’s rule, for the Graphical Lasso is:
-$\hat{\bm{\Theta}}=\operatorname*{argmin}_{\bm{\Theta}}-\log\det\bm{\Theta}+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{\Theta}\|_{1,\mathrm{off}}\Leftrightarrow\hat{\bm{\Theta}}^{-1}+\bm{S}\in\partial\|\hat{\bm{\Theta}}\|_{1,\mathrm{off}}$ (19)
-Now $\bm{U}\in\partial\|\hat{\bm{\Theta}}\|_{1,\mathrm{off}}$ if and only if
-  • $U_{ii}=0$ for $i\in[p]$
-  • $U_{ij}\in\lambda\partial|\Theta_{ij}|=\begin{cases}\lambda\operatorname{sign}(\Theta_{ij})$ if $\Theta_{ij}\neq 0\\ [-\lambda,\lambda]\text{ otherwise}\end{cases}$ for $i,j\in[p],i\neq j$.
-Hence
-[…]
-$\Leftrightarrow s_{ij}\in[-\lambda,\lambda]\quad\forall i\neq j$ (21)
-$\Leftrightarrow\lambda\geq\lambda_{\max}=\max_{i\neq j}|s_{ij}|$ (22)
+B.3 GLAD’s non-SPD outputs
+As Figure 8 shows, in the setting of Figure 4, GLAD’s best performing outputs $\bm{Z}$ are SPD in virtually $0\%$ if the case strongly sparse case with $p=100$.
+Figure 8: Percentage of SPD outputs among GLAD’s $\bm{Z}$ and $\bm{\Theta}$ outputs at convergence. As detailed in Section 4.1, $\bm{Z}$ is not guaranteed to be sparse.

### v4

@@
 Namely, additional structure such as sparsity can trivially be imposed on $\bm{\Theta}$ through $f$.
+Broadly speaking, constraints related to the values of the off-diagonal entries of $\bm{\Theta}$ and that can be imposed on the outputs of a neural network can be controlled directly.
 We next detail the computational cost of fully updating $\bm{\Theta}$ through one SpodNet layer.
@@
 $\begin{bmatrix}\bm{W}_{11}&\bm{w}_{12}\\ \bm{w}_{21}&w_{22}\end{bmatrix}$ $=\begin{bmatrix}\bm{\Theta}_{11}&\bm{\theta}_{12}\\ \bm{\theta}_{21}&\theta_{22}\end{bmatrix}^{-1}=\begin{bmatrix}[{\bm{\Theta}}_{11}]^{-1}+\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}&-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\\ \\ \left(-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\right)^{\top}&\frac{1}{\theta_{22}-\bm{{\theta}}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}\end{bmatrix}\,.$ (4)
-Identifying all blocks yields $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$ which can be computed in $\mathcal{O}(p^{2})$ if one has access to $\bm{W}$.
+Identifying all blocks (namely $g(\Theta)$ with $\theta_{22}-\theta_{21}[\Theta_{11}]^{-1}\theta_{12}$, $\bm{w}_{12}$ with $-\frac{[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}{\theta_{22}-\bm{\theta}_{21}[{\bm{\Theta}}_{11}]^{-1}\bm{\theta}_{12}}$ and $w_{22}$ with $\frac{1}{g(\bm{\Theta})}$) yields $[\bm{\Theta}_{11}]^{-1}=\bm{W}_{11}-\frac{1}{w_{22}}\bm{w}_{12}\bm{w}_{12}^{\top}$ which can be computed in $\mathcal{O}(p^{2})$ if one has access to $\bm{W}$.
 This can be achieved using Equation 3, and involves only operations in $\mathcal{O}(p^{2})$.
@@
 where $\|\bm{\Theta}\|_{1,\mathrm{off}}$ denote the off-diagonal $\ell_{1}$ norm of $\bm{\Theta}$, equal to $\sum_{i\neq j}|\Theta_{ij}|$.
+The data fidelity term $-\log\det(\bm{\Theta})+\langle\bm{S},\bm{\Theta}\rangle$ is the negative log-likelihood under Gaussian assumption, while the $\ell_{1}$ penalty enforces sparsity.
 It can be efficiently computed (Rolfs et al., 2012; Mazumder and Hastie, 2012; Hsieh et al., 2014) but suffers from the known limitations from the model-based approaches (Adler and Öktem, 2018; Arridge et al., 2019).
@@
 Its positivity is ensured by using an absolute value function as final nonlinearity.
+These input features are based on the intuition that the network should exploit information from the current state of the diagonal of $\bm{\Theta}$, the diagonal of $\bm{S}$ which is closely related to the diagonal of the GLasso estimator at convergence, and Schur’s complement which is added to the output of $g$.
 Improving training stability
@@
 Additional details regarding the experimental setups can be found in Appendix A.
-PyTorch implementations are provided with the submission and will be made public upon acceptance.
+PyTorch implementations can be found in our GitHub repository [^2] https://github.com/Perceptronium/SpodNet.
 5.1 Sparse precision matrix recovery
@@
 (2017).
-Figure 6: Graphs estimation on the Animals dataset with the GLasso, EGFM and our UBG model.
-The former and the latter yield cleaner and more interpretable clusters than the GLasso, but EGFM is specifically tailored for this task as opposed to our UBG model.
+Figure 6: Graph estimation on the Animals dataset with the GLasso, EGFM, GLAD and our UBG model.
+The last three yield cleaner and more interpretable clusters than the GLasso, but EGFM is specifically tailored for finding clusters as opposed to our UBG model or GLAD.
 Data & training
@@
 Baselines
-We compare our approach with the GLasso (with the regularization parameter selected by cross-validation), and the Elliptical Graphical Factor Model (EGFM) proposed by Hippert-Ferrer et al.
+We compare our approach with the GLasso (with the regularization parameter selected by cross-validation), GLAD (which is trained with the same data as our model), and the Elliptical Graphical Factor Model (EGFM) proposed by Hippert-Ferrer et al.
 (2023).
@@
 The results are illustrated in Figure 6, where the nodes of the graphs are partitioned using the Louvain algorithm (Blondel et al., 2008).
-Qualitatively, we observe that UBG produces a coherent graph, with outcomes comparable to those of GLasso and EGFM.
+Qualitatively, we observe that UBG produces a coherent graph, with outcomes comparable to those of the three baselines.
 Additionally, the graph learned by our UBG model exhibits a cleaner structure compared to GLasso, with a sparser representation and finer clusters, such as the grouping of (Chimp, Gorilla).
 To quantitatively assess the quality of the obtained graphs, we calculate the modularity $m$ of the partitions: higher modularity values indicate better separation of the graph into distinct subcomponents (Newman, 2006).
-We find that the graph produced by UBG achieves a higher modularity ($m=0.78$) than GLasso ($m=0.61$) and slightly lower than EGFM ($m=0.86$).
+We find that the graph produced by UBG achieves a higher modularity ($m=0.78$) than GLasso ($m=0.61$) and GLAD ($m=0.72$), and slightly lower than EGFM ($m=0.86$).
 It is important to note that the EGFM is specifically designed to obtain clustered graphs, whereas our UBG model has no such inductive bias.
@@
 We have shown how to leverage SpodNet to build neural architectures that outperform traditional methods in the context of sparse precision matrix estimation in various settings.
-Future directions of research include theoretical investigations in order to better understand the expressivity and derive convergence guarantees for our approach.
+Future research will be dedicated to improving the computational cost of our framework, theoretical understanding of the eigenvalues’ dynamics during training and of SpodNet’s expressivity, and deriving formal convergence guarantees.
 7 Reproducibility statement
@@
 The conducted experiments are described in Section 5 with additional details in Appendix A.
-PyTorch implementations are provided along with the submission.
-8 Acknowledgments
+PyTorch implementations can be found in our GitHub repository.
+8 Acknowledgements
 We would like to thank Badr Moufad (Ecole Polytechnique, France) for his help with the implementations, Paulo Gonçalves (Inria Lyon, France) for fruitful discussions as well as the Centre Blaise Pascal for computing ressources, which uses the SIDUS solution developed by Emmanuel Quemener (ENS Lyon, France) (Quemener and Corvellec, 2013).
@@
 As Figure 8 shows, in the setting of Figure 4, GLAD’s best performing outputs $\bm{Z}$ are SPD in virtually $0\%$ if the case strongly sparse case with $p=100$.
+Additionally, $\bm{Z}$ loses its sparsity if projected onto the SPD cone and becomes singular (since it is actually projected onto the closed cone of semi-definite matrices).
+Going the other way around and trying to sparsify $\bm{\Theta}$ instead, by thresholding its off-diagonal entries following the support found by $\bm{Z}$, breaks its positive-definiteness guarantee (Figure 9).
 Figure 8: Percentage of SPD outputs among GLAD’s $\bm{Z}$ and $\bm{\Theta}$ outputs at convergence. As detailed in Section 4.1, $\bm{Z}$ is not guaranteed to be sparse.
+Figure 9: First row: Projecting GLAD-$\bm{Z}$ onto the PSD cone. Second row: Thresholding GLAD-$\bm{\Theta}$ following the support of GLAD-$\bm{Z}$.